### v1

# Convergence to equilibrium for a degenerate three species reaction-diffusion system

Abstract.
In this work, we study a reaction-diffusion system for three species. The nonlinearity in the system comes from an underlying chemical reaction. Our main objective is to understand the long time behaviour of solutions to this reaction-diffusion system when there are degeneracies. More precisely, we treat cases when one of the diffusion coefficients vanishes while the other two diffusion coefficients stay positive. We prove convergence to equilibrium type results. In all our results, the constants appearing in the decay estimates are explicit.
1. Introduction
Reaction-diffusion equations are among the most widely used differential equations in applications. These equations govern the evolution (in time) of species concentrations at various spatial locations that are simultaneously diffusing and undergoing chemical reactions. We consider a reaction-diffusion model that concerns the three species $X_{1},X_{2}$ and $X_{3}$ undergoing the following reversible reaction:
$X_{1}+X_{2}\rightleftharpoons X_{3}$
The spatial domain is taken to be a bounded domain $\Omega\subset\mathbb{R}^{N}$ with $\mathrm{C}^{2+\nu}$ boundary with $\nu>0$. For the unknown $a,b,c:[0,T)\times\Omega\to\mathbb{R}$, representing the concentrations of 3 species $X_{1},X_{2},X_{3}$ respectively, the reaction-diffusion model is given by the following system of partial differential equations:
(1) $\partial_{t}a-d_{a}\Delta a&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}b-d_{b}\Delta b&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}c-d_{c}\Delta c&=ab-c\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ n\cdot\nabla_{x}a=n\cdot\nabla_{x}b&=n\cdot\nabla_{x}c=0\quad\mbox{ on }(0,T)\times\partial\Omega,\\ a(0,\cdot)=a_{0};\,b(0,\cdot)=b_{0};\,&c(0,\cdot)=c_{0}\qquad\qquad\mbox{ in }\Omega.\end{aligned}\right.$
Here $n(x)$ denotes the outward unit normal to $\Omega$ at the point $x\in\partial\Omega$. The initial data $a_{0},b_{0},c_{0}$ are taken to be smooth and the diffusion coefficients $d_{a},d_{b},d_{c}$ are taken to be nonnegative. When the diffusion coefficients are strictly positive (referred to, now onwards, as the non-degenerate setting), it is well-known that a unique strictly positive global-in-time $\mathrm{C}^{\infty}$ solution exists for the above model [Pie04, Pie10], where the positiveness comes from the quasi-positive nature of the rate function [Ama85, QS19, Rot06]. The large time behaviour of solutions to (1) in the non-degenerate setting was addressed in [DF06] using the method of entropy. In this method, a Lyapunov functional (termed entropy) is found for the evolution equation. The negative of the time derivative of this entropy functional is referred to as the entropy dissipation functional. The entropy dissipation functional is then related back to the relative entropy via a functional inequality. This will then be followed by a Grönwall type argument to deduce convergence of relative entropy to zero. A Cziszár-Kullback-Pinsker type inequality that relates relative entropy and the $\mathrm{L}^{1}$-norm helps the authors in [DF06] to prove the convergence to equilibrium in the $\mathrm{L}^{1}(\Omega)$-norm and that the decay is exponentially fast in time. In the proof of [DF06], uniform boundedness of the solution $(a,b,c)$ to (1) plays an important role. This uniform bound is available in the non-degenerate case.
In this paper, we discuss the large time behaviour of solutions to couple of degenerate cases (i.e when one of the species stops diffusing) of the three species model (1). Our work is heavily inspired by [EMT20] which dealt with a four species degenerate model where one of the species stops diffusing. The authors in [EMT20] demonstrated a so-called indirect diffusion effect wherein an effective diffusion is felt by the non-diffusive species, thanks to interplay between the diffusion from diffusive species and the underlying reversible reaction. Our work demonstrates that a similar indirect diffusion effect is present in the above three species model in the presence of degeneracies. The two degenerate reaction-diffusion systems that we study in this article are:
(2) $\partial_{t}a-d_{a}\Delta a&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}b&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}c-d_{c}\Delta c&=ab-c\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ n\cdot\nabla_{x}a=n\cdot\nabla_{x}b&=n\cdot\nabla_{x}c=0\quad\mbox{ on }(0,T)\times\partial\Omega\\ a(0,\cdot)=a_{0};\,b(0,\cdot)=b_{0};\,&c(0,\cdot)=c_{0}\qquad\qquad\mbox{ in }\Omega.\end{aligned}\right.$
and
(3) $\partial_{t}a-d_{a}\Delta a&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}b-d_{b}\Delta b&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}c&=ab-c\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ n\cdot\nabla_{x}a=n\cdot\nabla_{x}b&=n\cdot\nabla_{x}c=0\quad\mbox{ on }(0,T)\times\partial\Omega\\ a(0,\cdot)=a_{0};\,b(0,\cdot)=b_{0};\,&c(0,\cdot)=c_{0}\qquad\qquad\mbox{ in }\Omega.\end{aligned}\right.$
Here, the initial data are assumed to be smooth and strictly positive. The diffusion coefficients $d_{a},d_{c}$ in (2) and the diffusion coefficients $d_{a},d_{b}$ in (3) are assumed to be strictly positive. The existence of a smooth positive solution to (2) was proved in [DF15, Theorem 3.1] for all smooth initial non-negative data. The existence of a smooth positive solution to (3) was proved in [DF15, Theorem 3.2] for all smooth initial non-negativce data if the dimension $N\leq 3$. The authors in [DF15], however, prove the existence of a weak global-in-time positive solution to (3) in any arbitrary dimension. The process of establishing various estimates in the article [EMT20] has greatly motivated us to establish similar kind of estimates in dimension larger than three for the degenerate case (2) corresponding to $d_{b}=0$. We further use various Neumann Green’s function results from [FMT19, Rot06, Mor83, CK15], to get various estimates specifically for dimensions less than four.
Observe that both the models (2) and (3) satisfy the following mass conservation properties:
(4) $0<M_{1}|\Omega|=:\int_{\Omega}a(t,x)+c(t,x)=\int_{\Omega}a(0,x)+c(0,x)$
(5) $0<M_{2}|\Omega|=:\int_{\Omega}b(t,x)+c(t,x)=\int_{\Omega}b(0,x)+c(0,x)$
An equilibrium state $(a_{\infty},b_{\infty},c_{\infty})$ associated with these models should also satisfy these conservation properties. Hence we should have
(6) $a_{\infty}+c_{\infty}=M_{1},\qquad b_{\infty}+c_{\infty}=M_{2}$
Moreover, at equilibrium the rate function should vanish, i.e.
(7) $c_{\infty}=a_{\infty}b_{\infty}$
The relations (6) and (7) put together leads to a quadratic equation for $c_{\infty}$ whose only admissible non-negative solution is
(8) $c_{\infty}=\frac{1}{2}(1+M_{1}+M_{2})-\frac{1}{2}\sqrt{(1+M_{1}+M_{2})^{2}-4M_{1}M_{2}}.$
From now onwards, we will be considering this unique homogeneous equilibrium state $(a_{\infty},b_{\infty},c_{\infty})$.
Some Notation:
  • For any function $f:\Omega\to\mathbb{R}$ we denote the average of $f$ as
    $\overline{f}=\frac{1}{|\Omega|}\int_{\Omega}f(x)\,{\rm d}x.$
  • $\Omega_{\tau,T}=(\tau,T)\times\Omega$ and $\Omega_{t}=(0,t)\times\Omega$
  • For $a,b,c$ non-negative, we define $A:=\sqrt{a};\ B:=\sqrt{b};\ C:=\sqrt{c}$. The deviations from their averages are denoted by
    $\delta_{A}:=A-\overline{A};\quad\delta_{B}:=B-\overline{B};\quad\delta_{C}:=C-\overline{C}.$
We consider the following entropy functional associated with (2) and (3):
(9) $E(a,b,c)=\int_{\Omega}(a(\ln a-1)+1)+(b(\ln b-1)+1)+(c(\ln c-1)+1).$
While studying (2) in dimensions $N\geq 4$, we have been able to arrive at large time asymptotics of the solution only under certain closeness assumption on the non-zero diffusion coefficients. The precise assumption is the following:
Assumption 1.
The non-zero diffusion coefficients $d_{a}$ and $d_{c}$ are said to satisfy the closeness assumption if
(10) $|{d_{a}-d_{c}}|<\frac{2}{C^{PRC}_{{\frac{d_{a}+d_{c}}{2}},p^{{}^{\prime}}}}\qquad\text{ and }\qquad\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}<\frac{1}{C_{SOR}(\Omega,N,p^{{}^{\prime}})},$
where the constants $C^{PRC}_{{\frac{d_{a}+d_{c}}{2}},p^{{}^{\prime}}}$ and $C_{SOR}(\Omega,N,p^{{}^{\prime}})$, are the parabolic regularity constant (see Theorem 12 in the Appendix) and the second order regularity constant (see Theorem 11 in the Appendix), respectively.
Now we are going to state our two main results.
Theorem 1.
For $N\geq 4$, let $(a,b,c)$ be the solution to the degenerate system (2). Let $(a_{\infty},b_{\infty},c_{\infty})$ be the associated equilibrium state given by (6)-(8). Let the nonzero diffusion coefficients $d_{a},d_{c}$ satisfy the closeness condition (10). Then, for any given positive $\varepsilon\ll 1$, there exists a time $T_{\varepsilon}$ and two positive constants $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ such that for $t\geq T_{\varepsilon}$, we have
$\frac{1}{2M_{1}}\|a-a_{\infty}\|_{L^{1}(\Omega)}^{2}+\frac{1}{2M_{2}}\|b-b_{\infty}\|_{L^{1}(\Omega)}^{2}+\frac{1}{(M_{1}+M_{2})}\|c-c_{\infty}\|_{L^{1}(\Omega)}^{2}\leq\frac{(9+2\sqrt{2})}{(3+2\sqrt{2})|\Omega|}\mathcal{S}_{1}e^{-\mathcal{S}_{2}(1+t)^{\frac{1-\epsilon}{N-1}}}.$
For $N<4$, let $(a,b,c)$ be the solution to the degenerate system (2). Let $(a_{\infty},b_{\infty},c_{\infty})$ be the associated equilibrium state given by (6)-(8). Then, for any given positive $\varepsilon\ll 1$, there exists a time $T_{\varepsilon}$ and two positive constants $\mathcal{S}_{3}$ and $\mathcal{S}_{4}$ such that for $t\geq T_{\varepsilon}$, we have
$\frac{1}{2M_{1}}\|a-a_{\infty}\|_{L^{1}(\Omega)}^{2}+\frac{1}{2M_{2}}\|b-b_{\infty}\|_{L^{1}(\Omega)}^{2}+\frac{1}{(M_{1}+M_{2})}\|c-c_{\infty}\|_{L^{1}(\Omega)}^{2}\leq\frac{(9+2\sqrt{2})}{(3+2\sqrt{2})|\Omega|}\mathcal{S}_{3}\,e^{-\mathcal{S}_{4}(1+t)^{\frac{1-\epsilon}{6}}}.$
Theorem 2.
Let $N\leq 3$ and let $(a,b,c)$ be the solution to the degenerate system (3). Let $(a_{\infty},b_{\infty},c_{\infty})$ be the associated equilibrium state given by (6)-(8). Then, for any given positive $\varepsilon\ll 1$, there exists a time $T_{\varepsilon}$ and two positive constants $\mathcal{S}_{5}$ and $\mathcal{S}_{6}$ such that for $t\geq T_{\varepsilon}$, we have
$\frac{1}{2M_{1}}\|a-a_{\infty}\|_{L^{1}(\Omega)}^{2}+\frac{1}{2M_{2}}\|b-b_{\infty}\|_{L^{1}(\Omega)}^{2}+\frac{1}{(M_{1}+M_{2})}\|c-c_{\infty}\|_{L^{1}(\Omega)}^{2}\leq\frac{(9+2\sqrt{2})}{(3+2\sqrt{2})|\Omega|}\mathcal{S}_{5}\,e^{-\mathcal{S}_{6}(1+t)^{\frac{2-\epsilon}{3}}}.$
Now based on solutions of two degenerate systems we define two different entropy dissipation functional which will be related to the entropy functional by a sign of time derivative.
We define the following positive entropy dissipation functionals for $d_{b}=0$ and $d_{c}=0$:
(11) $\text{For}\ d_{b}&=0\\ D(a,b,c)=&d_{a}\int_{\Omega}\frac{|\nabla a|^{2}}{a}\ +d_{c}\int_{\Omega}\frac{|\nabla c|^{2}}{c}\ +\int_{\Omega}(ab-c)\ln\Big{(}\frac{ab}{c}\Big{)}.\\ \text{For}\ d_{c}&=0\\ D(a,b,c)=&d_{a}\int_{\Omega}\frac{|\nabla a|^{2}}{a}\ +d_{b}\int_{\Omega}\frac{|\nabla b|^{2}}{b}\ +\int_{\Omega}(ab-c)\ln\Big{(}\frac{ab}{c}\Big{)}.\end{aligned}\right.$
Using classical Poincaré inequality, taking $C(\Omega)$ as the Poincaré constant of the domain, and on the last term application of the following algebraic inequality
$(x-y)(\ln{x}-\ln{y})\geq 4(\sqrt{x}-\sqrt{y})^{2};\ \ \forall x,y\geq 0,\ \text{yields:}$:
(12) {$D(a,b,c)\geq&\frac{d_{a}}{C(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{d_{c}}{C(\Omega)}\|\delta_{C}\|^{2}_{\mathrm{L}^{2}(\Omega)}+4\|AB-C\|^{2}_{L^{2}(\Omega)}\qquad\mbox{ for }\,d_{b}=0,\\[5.69046pt] D(a,b,c)\geq&\frac{d_{a}}{C(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{d_{b}}{C(\Omega)}\|\delta_{B}\|^{2}_{\mathrm{L}^{2}(\Omega)}+4\|AB-C\|^{2}_{L^{2}(\Omega)}\qquad\mbox{ for }\,d_{c}=0.\end{aligned}\right.$
An application of Poincaré-Writinger inequality (see Lemma 7 in the Appendix) for the dimension  $N\geq 4$ gives us for $d_{b}=0$,
$D(a,b,c)\geq\frac{d_{a}}{P(\Omega)}\|\delta_{A}\|^{2}_{L^{\frac{2N}{N-2}}(\Omega)}+\frac{d_{c}}{P(\Omega)}\|\delta_{C}\|^{2}_{L^{\frac{2N}{N-2}}(\Omega)}+\|AB-C\|^{2}_{L^{2}(\Omega)}$
and the same inequality for dimension $N=1,2,3$, gives us for $d_{b}=0$,
$D(a,b,c)\geq\frac{4d_{a}}{P(\Omega)}\|\delta_{A}\|^{2}_{L^{6}(\Omega)}+\frac{4d_{c}}{P(\Omega)}\|\delta_{C}\|^{2}_{L^{6}(\Omega)}+\|AB-C\|^{2}_{L^{2}(\Omega)}$
where $P(\Omega)$ is the Poincaré-Writinger constant (see Lemma 7 in the Appendix).
A similar application of Poincaré-Writinger inequality for $N\geq 4$ gives us for $d_{c}=0$,
$D(a,b,c)\geq\frac{4d_{a}}{P(\Omega)}\|\delta_{A}\|^{2}_{L^{\frac{2N}{N-2}}(\Omega)}+\frac{4d_{b}}{P(\Omega)}\|\delta_{B}\|^{2}_{L^{\frac{2N}{N-2}}(\Omega)}+\|AB-C\|^{2}_{L^{2}(\Omega)}$
and the same inequality for the dimension $N=1,2,3$, gives us for $d_{c}=0$,
$D(a,b,c)\geq\frac{4d_{a}}{P(\Omega)}\|\delta_{A}\|^{2}_{L^{6}(\Omega)}+\frac{4d_{b}}{P(\Omega)}\|\delta_{B}\|^{2}_{L^{6}(\Omega)}+\|AB-C\|^{2}_{L^{2}(\Omega)}$
where $P(\Omega)$ is the Poincaré-Writinger constant (see Lemma 7 in the Appendix)
Furthermore for both the cases we have the following relation between the two functional by the time derivative i.e
$\frac{d}{dt}(E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty}))=-D(a,b,c)$
Furthermore consider the function $\Gamma(x,y)=\frac{x\ln{\frac{x}{y}}-x+y}{(\sqrt{x}-\sqrt{y})^{2}}$ ; we see that we can find a Constant $C_{\Gamma}>0$; such that $\Gamma(x,y)\leq C_{\Gamma}\max\{1,\ln{\frac{x}{y}}\}$. We are going to represent difference between entropy by this function. We can write:
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})$
$=\int_{\Omega}\Gamma(a,a_{\infty})(A-A_{\infty})^{2}$ $+\int_{\Omega}\Gamma(b,b_{\infty})(B-B_{\infty})^{2}+\int_{\Omega}\Gamma(c,c_{\infty})(C-C_{\infty})^{2}$
$\leq C_{\Gamma}\max\big{\{}1,\ln{(\|a\|_{L^{\infty}(\Omega)}+1)}+$ $|\ln{a_{\infty}}|\big{\}}\|A-A_{\infty}\|_{L^{2}(\Omega)}^{2}$
$+C_{\Gamma}\max\big{\{}1,\ln{(\|b\|_{L^{\infty}(\Omega)}+1)}$ $+|\ln{b_{\infty}}|\big{\}}\|B-B_{\infty}\|_{L^{2}(\Omega)}^{2}$
$+C_{\Gamma}\max\big{\{}1,\ln{(\|c\|_{L^{\infty}(\Omega)}+1)}$ $+|\ln{c_{\infty}}|\big{\}}\|C-C_{\infty}\|_{L^{2}(\Omega)}^{2}$
Now for degeneracy $d_{b}=0$, we derive in Lemma 4 and for degeneracy $d_{c}=0$, we derive in Proposition 8 that the growth of the species varies polynomially with time, which leads us to the following relation: for some positive constant $C_{1}>0$, and time sufficiently large time i.e $t\geq T_{\epsilon}$, where $\epsilon>0$,
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})$
$\leq$ $C_{1}(1+t)^{\epsilon}\Big{(}\|A-A_{\infty}\|_{L^{2}(\Omega)}^{2}+\|B-B_{\infty}\|_{L^{2}(\Omega)}^{2}+\|C-C_{\infty}\|_{L^{2}(\Omega)}^{2}\Big{)}$
Now we have from [DF06]; $\exists C_{EB}>0$, depending only on the domain and equilibrium point such that:
$\|\sqrt{\overline{A^{2}}}-A_{\infty}$ $\|_{L^{2}(\Omega)}^{2}+\|\sqrt{\overline{B^{2}}}-B_{\infty}\|_{L^{2}(\Omega)}^{2}+\|\sqrt{\overline{C^{2}}}-C_{\infty}\|_{L^{2}(\Omega)}^{2}$
$\leq$ $C_{EB}\big{(}\|\delta_{A}\|_{L^{2}(\Omega)}^{2}+\|\delta_{B}\|_{L^{2}(\Omega)}^{2}+\|\delta_{C}\|_{L^{2}(\Omega)}^{2}+\|C-AB\|_{L^{2}(\Omega)}^{2}\big{)}$
This indicates that we will able to apply Grónwall inequality if we somehow link entropy dissipation with the missing term $\|\delta_{B}\|^{2}_{L^{2}(\Omega)}$ for $d_{b}=0$ case and similarly with the missing term $\|\delta_{C}\|^{2}_{L^{2}(\Omega)}$ for $d_{c}=0$ case. So to obtain exponential decay in convergence type of results our goal is to relate entropy dissipation with the missing term.
We will relate the missing term with entropy dissipation in Proposition 4 for $d_{b}=0$ case in dimension $N\geq 4$, and proved the same for $d_{b}=0$ case in dimension $N=1,2,3$ in Proposition 5 in the following way:
[…]
To prove this we require an estimate of $L^{\frac{N}{2}}(\Omega)$ norm of $b$ for $N\geq 4$ and an estimate of $L^{\frac{3}{2}}(\Omega)$ norm of $b$ for $N=1,2,3$. These estimates are the Key Estimates of our article. It turns out our Key estimates are the following:
  • For $N\geq 4$   $\|b\|_{L^{\frac{N}{2}}(\Omega)}\leq\hat{K}(1+t)^{\frac{N-2}{N-1}}$
  • For $N=1,2,3;\ \ \|b\|_{L^{\frac{3}{2}}(\Omega)}\leq\hat{K}(1+t)^{\frac{5}{6}}$
[…]
$\|a\|_{L^{p}(\Omega_{\tau,\tau+1})}+\|c\|_{L^{p}(\Omega_{\tau,\tau+1})}\leq C_{0}\qquad\forall\tau\geq 0$
This above result is proved in Lemma 1, and this will help us to prove our key estimate for dimension $N\geq 4$. Similarly we will able to prove time dependent evolution of space integral of diffusive species for degenerate $d_{b}=0$ case in dimension $N=1,2,3$, in Proposition 3 and Lemma 3. The relation for dimension $N=1,2,3$ for $d_{b}=0$ case is described below:for some natural number $\mu_{C}>0$,
$\|a\|_{L^{p}(\Omega_{\tau,\tau+1})}+\|c\|_{L^{p}(\Omega_{\tau,\tau+1})}\leq K_{c}(1+\tau)^{\mu_{C}}\qquad\forall\tau\geq 0$
In this article our calculations are inspired from the article [EMT20].
For $d_{c}=0$, we will relate entropy dissipation with $\delta_{C}$. In this article for $d_{c}=0$ we consider the dimension $N\leq 4$. Although we will establish our result only for dimension $N=3$, for dimension $N=1,2,4$ all the calculations are similar. In Proposition 9 We establish the following relations between entropy dissipation and $\delta_{C}$:
$D(a,b,c)\geq\hat{C}(1+t)^{-\frac{1}{3}}\big{(}\|A-\overline{A}\|_{L^{2}(\Omega)}^{2}+\|B-\overline{B}\|_{L^{2}(\Omega)}^{2}+\|C-\overline{C}\|_{L^{2}(\Omega)}^{2}\big{)}$
To prove this we need a particular integral estimate of species $a$, which we will establish in Lemma 6. The particular relation is the following:
$\|a\|_{L^{\frac{3}{2}}(\Omega)}\leq\tilde{C}(1+t)^{\frac{1}{3}};\qquad\forall t\geq 0.$
This is the key estimate for the degenerate case $d_{c}=0$; to establish this we will use the following estimate (see Lemma 5): $\exists C_{\Omega}>0$ independent of time such that;
$\int_{0}^{T}\int_{\Omega}a^{2}\leq C_{\Omega}(1+T);\ \ \int_{0}^{T}\int_{\Omega}b^{2}\leq C_{\Omega}(1+T)\qquad\forall T\geq 0$
These results along with Grónwall inequality will help us to show entropy functional for both degeneracy cases $d_{b}=0$ and $d_{c}=0$ decays sub-exponentially with time. On the other hand an application of Cziszár-Kullback-Pinsker Inequality provides us the following lower bound of entropy functional; [DF06]
$E(a,b,c)$ $-E(a_{\infty},b_{\infty},c_{\infty})\geq\frac{(3+2\sqrt{2})|\Omega|}{2M_{1}(9+2\sqrt{2})}\|a-a_{\infty}\|_{L^{1}(\Omega)}^{2}$
$+$ $\frac{(3+2\sqrt{2})|\Omega|}{2M_{2}(9+2\sqrt{2})}\|b-b_{\infty}\|_{L^{1}(\Omega)}^{2}+\frac{(3+2\sqrt{2})|\Omega|}{(M_{1}+M_{2})(9+2\sqrt{2})}\|c-c_{\infty}\|_{L^{1}(\Omega)}^{2}$
Combining all these we arrive at the main results of our article i.e Theorem 1 and Theorem 2.
2. The case of $d_{b}=0$
The idea is to connect entropy dissipation with the missing $\delta_{B}$ term in (12) so that we can apply Grönwall inequality to have a sub-exponential decay. We begin by proving an uniform integrability estimate for $a$ and $c$ in a parabolic cylinder of unit height. A similar estimate was obtained for a degenerate four species model in [EMT20, Lemma 3.12, p.4343].
Lemma 1.
Let $p>N\geq 4$ and let $p^{\prime}$ be its Hölder conjugate. Let $(a,b,c)$ be the solution to the degenerate system (2) and let the nonzero diffusion coefficients $d_{a},d_{c}$ satisfy the closeness condition (10). Then there exists a $C_{0}>0$, depending only on the initial condition and the dimension $N$, such that
$\|a\|_{\mathrm{L}^{p}(\Omega_{\tau,\tau+1})}+\|c\|_{\mathrm{L}^{p}(\Omega_{\tau,\tau+1})}\leq C_{0}\qquad\forall\tau>0.$
Proof.
Define $d:=\frac{d_{a}+d_{c}}{2}$. We can rewrite the equations corresponding to concentrations $a$ and $c$ in (2) as
[…]
Let $\phi:[0,\infty)\rightarrow[0,1]$ be a smooth function such that $\phi(0)=0$ and
$\phi(x)$ $=1\qquad\mbox{ for }x\in[1,\infty)$
$0\leq\phi^{\prime}(x)$ $\leq M\qquad\mbox{ for }x\in[0,\infty)$
for some constant $M>0$. For an arbitrary $\tau>0$, consider $\phi_{\tau}:[\tau,\infty)\to[0,1]$ defined as $\phi_{\tau}(s):=\phi(s-\tau)$ for $s\in[\tau,\infty)$. Then, the product $\phi_{\tau}(t)a(t,x)$ satisfies
$\partial_{t}\left(\phi_{\tau}a\right)-d\Delta\left(\phi_{\tau}a\right)&=a\partial_{t}\phi_{\tau}+\phi_{\tau}(c-ab+(d_{a}-d)\Delta a)\qquad\mbox{ in }\Omega_{\tau,\tau+2}\\ \phi_{\tau}\nabla a\cdot n&=0\qquad\qquad\qquad\qquad\ \qquad\qquad\qquad\ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}(\tau)a(\tau,x)&=0\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ \ \ \mbox{ in }\Omega\end{aligned}\right.$
Making the change of variable $t_{1}=t-\tau$ in the above equation yields
(14) $\partial_{t_{1}}(\phi_{\tau}(t_{1}+\tau)a(t_{1}+\tau,x))&-d\Delta(\phi_{\tau}(t_{1}+\tau)a(t_{1}+\tau,x))\\ =a(t_{1}+\tau,x)\partial_{t_{1}}\phi_{\tau}(t_{1}&+\tau)+\phi_{\tau}(t_{1}+\tau)(c-ab+(d_{a}-d)\Delta a)\qquad\mbox{ in }\Omega_{0,2}\\ \phi_{\tau}(t_{1}+\tau)\nabla a(t_{1}+\tau,x)\cdot n&=0\qquad\qquad\quad\qquad\qquad\qquad\qquad\qquad\qquad\mbox{ on }\partial\Omega_{0,2}\\ \phi_{\tau}(0+\tau)a(0+\tau,x)&=0\qquad\qquad\qquad\qquad\quad\qquad\qquad\qquad\qquad\mbox{ in }\Omega\end{aligned}\right.$
Similarly, the concentration $c$ satisfies the following boundary value problem:
(15) $\partial_{t_{1}}(\phi_{\tau}(t_{1}+\tau)c(t_{1}+\tau,x))&-d\Delta(\phi_{\tau}(t_{1}+\tau)c(t_{1}+\tau,x))\\ =c(t_{1}+\tau,x)\partial_{t_{1}}\phi_{\tau}(t_{1}&+\tau)+\phi_{\tau}(t_{1}+\tau)(ab-c+(d_{c}-d)\Delta c)\qquad\mbox{ in }\Omega_{0,2}\\ \phi_{\tau}(t_{1}+\tau)\nabla c(t_{1}+\tau,x)\cdot n&=0\qquad\quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\mbox{ on }\partial\Omega_{0,2}\\ \phi_{\tau}(0+\tau)c(0+\tau,x)&=0\qquad\qquad\qquad\qquad\quad\qquad\qquad\qquad\qquad\mbox{ in }\Omega\end{aligned}\right.$
Let $G_{d}$ denotes the Green’s function associated with the operator $\partial_{t}-d\Delta$ with Neumann boundary condition. Then, we can express the solutions to (14) and (15) as follows:
(16) $\phi_{\tau}(t_{1}+\tau)a(t_{1}+\tau,x)$ $=\int_{0}^{t_{1}}\int_{\Omega}G_{d}(t_{1},s,x,y)a(s+\tau,y)\partial_{s}\phi_{\tau}(s+\tau)\,{\rm d}y\,{\rm d}s$
$+\int_{0}^{t_{1}}\int_{\Omega}G_{d}(t_{1},s,x,y)\phi_{\tau}(s+\tau)(c-ab+(d_{a}-d)\Delta a)(s,y)\,{\rm d}y\,{\rm d}s$
and
(17) $\phi_{\tau}(t_{1}+\tau)c(t_{1}+\tau,x)$ $=\int_{0}^{t_{1}}\int_{\Omega}G_{d}(t_{1},s,x,y)c(s+\tau,y)\partial_{s}\phi_{\tau}(s+\tau)\,{\rm d}y\,{\rm d}s$
$+\int_{0}^{t_{1}}\int_{\Omega}G_{d}(t_{1},s,x,y)\phi_{\tau}(s+\tau)(ab-c+(d_{c}-d)\Delta c)(s,y)\,{\rm d}y\,{\rm d}s$
Let us fix a non-negative $\theta\in\mathrm{L}^{p^{\prime}}(\Omega_{0,2})$. Let $\psi$ be the solution to
(18) $\partial_{t_{1}}\psi(t_{1},x)+d\Delta\psi(t_{1},x)&=-\theta(t_{1},x)\qquad\mbox{ for }(t_{1},x)\in\Omega_{0,2}\\ \nabla\psi(t_{1},x)\cdot n&=0\qquad\ \ \qquad\mbox{ for }(t_{1},x)\in(0,2)\times\partial\Omega\\ \psi(2,x)&=0\qquad\qquad\mbox{ for }x\in\Omega\end{aligned}\right.$
Applying the second order regularity estimate (see Theorem 11 in Appendix A for further details) to the above equation yields
(19) $\|\Delta\psi\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}\leq\frac{2C_{SOR}(\Omega,N,p^{\prime})}{d_{a}+d_{c}}\|\theta\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}$
Multiplying the expression (16) by $\theta$ and integrating over time and space yields
$\int_{0}^{2}\int_{\Omega}$ $\phi_{\tau}(t_{1}+\tau)a(t_{1}+\tau,x)\theta(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
$=\int_{0}^{2}\int_{\Omega}\left(\int_{0}^{t_{1}}\int_{\Omega}G_{d}(t_{1},s,x,y)a(s+\tau,y)\partial_{s}\phi_{\tau}(s+\tau)\,{\rm d}y\,{\rm d}s\right)\theta(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
$+\int_{0}^{2}\int_{\Omega}\left(\int_{0}^{t_{1}}\int_{\Omega}G_{d}(t_{1},s,x,y)\phi_{\tau}(s+\tau)(c-ab+(d_{a}-d)\Delta a)(s,y)\,{\rm d}y\,{\rm d}s\right)\theta(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
Substituting for $\theta$ in terms of $\psi$, using (18), in the second term on the right hand side of the above expression followed by integration by parts yields
$\int_{0}^{2}\int_{\Omega}$ $\phi_{\tau}(t_{1}+\tau)a(t_{1}+\tau,x)\theta(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
$=\int_{0}^{2}\int_{\Omega}\left(\int_{0}^{t_{1}}\int_{\Omega}G_{d}(t_{1},s,x,y)a(s+\tau,y)\partial_{s}\phi_{\tau}(s+\tau)\,{\rm d}y\,{\rm d}s\right)\theta(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
$+\int_{0}^{2}\int_{\Omega}\left(\int_{0}^{t_{1}}\int_{\Omega}\left(\partial_{t_{1}}-d\Delta\right)G_{d}(t_{1},s,x,y)\left(\phi_{\tau}(s+\tau)(c-ab+(d_{a}-d)\Delta a)(s,y)\right)\,{\rm d}y\,{\rm d}s\right)\psi(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
Using the property of the Green’s function, we get
$\int_{0}^{2}\int_{\Omega}$ $\phi_{\tau}(t_{1}+\tau)a(t_{1}+\tau,x)\theta(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
$=\int_{0}^{2}\int_{\Omega}\left(\int_{0}^{t_{1}}\int_{\Omega}G_{d}(t_{1},s,x,y)a(s+\tau,y)\partial_{s}\phi_{\tau}(s+\tau)\,{\rm d}y\,{\rm d}s\right)\theta(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
$+\int_{0}^{2}\int_{\Omega}\phi_{\tau}(t_{1}+\tau)(c-ab+(d_{a}-d)\Delta a)(t_{1},x)\psi(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
A further integration by parts in the second term on the right hand side of the above expression yields
(20) $\int_{0}^{2}\int_{\Omega}$ $\phi_{\tau}(t_{1}+\tau)a(t_{1}+\tau,x)\theta(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
$\leq\left\|\int_{0}^{t_{1}}\int_{\Omega}G_{d}(t_{1},s,x,y)a(s+\tau,y)\partial_{s}\phi_{\tau}(s+\tau)\,{\rm d}y\,{\rm d}s\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}\left\|\theta\right\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}$
$+\int_{0}^{2}\int_{\Omega}\phi_{\tau}(t_{1}+\tau)(\left(c-ab\right)\psi+(d_{a}-d)a\Delta\psi)(t_{1},x)\,{\rm d}x\,{\rm d}t_{1},$
where we have also used Hölder inequality to bound the first term on the right hand side.
The following Green’s function estimate is available from [Mor83][CK15]: there exists a constant $K_{1}>0$, depending only on the domain, such that
$0\leq G_{d}(t_{1},s,x,y)\leq\frac{K_{1}}{(4\pi(t_{1}-s))^{\frac{N}{2}}}e^{-\kappa\frac{\|x-y\|^{2}}{(t_{1}-s)}}=:g_{d}(t_{1}-s,x-y),$
for some constant $\kappa>0$ depending only on $\Omega$ and the diffusion coefficient $d$. Note that
$\left\|g_{d}\right\|_{L^{1+\frac{2}{N}}(\Omega_{0,2})}\leq K_{2}$
for some constant $K_{2}>0$. Observe that there exists a $q<p$ such that
$1+\frac{1}{p}=\frac{1}{1+\frac{2}{N}}+\frac{1}{q}.$
Hence, applying the Young’s convolution inequality we obtain
$\left\|\int_{0}^{t_{1}}\int_{\Omega}G_{d}(t_{1},s,x,y)a(s+\tau,y)\partial_{s}\phi_{\tau}(s+\tau)\,{\rm d}y\,{\rm d}s\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}$ $\leq\left\|g_{d}\right\|_{\mathrm{L}^{1+\frac{2}{n}}(\Omega_{0,2})}\left\|a(\cdot+\tau,\cdot)\partial_{s}\phi_{\tau}(\cdot+\tau)\right\|_{\mathrm{L}^{q}(\Omega_{0,2})}$
$\leq K_{2}\left\|a(\cdot+\tau,\cdot)\partial_{s}\phi_{\tau}(\cdot+\tau)\right\|_{\mathrm{L}^{q}(\Omega_{0,2})}$
Using this estimate and the estimate of $\Delta\psi$ from (19) in (20), we arrive at
(21) $\int_{0}^{2}\int_{\Omega}$ $\phi_{\tau}(t_{1}+\tau)a(t_{1}+\tau,x)\theta(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
$\leq K_{2}\left\|a(\cdot+\tau,\cdot)\partial_{s}\phi_{\tau}(\cdot+\tau)\right\|_{\mathrm{L}^{q}(\Omega_{0,2})}\left\|\theta\right\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}+\int_{0}^{2}\int_{\Omega}\phi_{\tau}(t_{1}+\tau)\left(\left(c-ab\right)\psi\right)(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
$+C_{SOR}(\Omega,N,p^{\prime})\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}\left\|\phi_{\tau}(\cdot+\tau)a(\cdot+\tau,\cdot)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}\|\theta\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}$
Performing a similar set of computations on the equation (17) for $\phi_{\tau}c$ yields
(22) $\int_{0}^{2}\int_{\Omega}$ $\phi_{\tau}(t_{1}+\tau)c(t_{1}+\tau,x)\theta(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
$\leq K_{2}\left\|c(\cdot+\tau,\cdot)\partial_{s}\phi_{\tau}(\cdot+\tau)\right\|_{\mathrm{L}^{q}(\Omega_{0,2})}\left\|\theta\right\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}+\int_{0}^{2}\int_{\Omega}\phi_{\tau}(t_{1}+\tau)\left(\left(ab-c\right)\psi\right)(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
$+C_{SOR}(\Omega,N,p^{\prime})\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}\left\|\phi_{\tau}(\cdot+\tau)c(\cdot+\tau,\cdot)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}\|\theta\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}$
Adding the inequalities (21) and (22) and using positivity of $a$ and $c$ yields
$\int_{0}^{2}\int_{\Omega}$ $\phi_{\tau}(t_{1}+\tau)\left(a+c\right)(t_{1}+\tau,x)\theta(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
$\leq K_{2}\left\|\left(a+c\right)(\cdot+\tau,\cdot)\partial_{s}\phi_{\tau}(\cdot+\tau)\right\|_{\mathrm{L}^{q}(\Omega_{0,2})}\left\|\theta\right\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}$
$+C_{SOR}(\Omega,N,p^{\prime})\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}\left\|\phi_{\tau}(\cdot+\tau)\left(a+c\right)(\cdot+\tau,\cdot)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}\|\theta\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}$
As $1\leq q<p$, there exists $\alpha\in(0,1]$ such that
$\frac{1}{q}=\frac{1-\alpha}{p}+\frac{\alpha}{1}.$
Hence by interpolation, we have
$\|f\|_{\mathrm{L}^{q}}\leq\|f\|_{\mathrm{L}^{p}}^{1-\alpha}\|f\|_{\mathrm{L}^{1}}^{\alpha}\qquad\forall f\in\mathrm{L}^{q}\cap\mathrm{L}^{1}.$
Using the above interpolation inequality and a duality argument, we obtain
(23) $\left\|\phi_{\tau}(\cdot+\tau)(a+c)(\cdot+\tau,\cdot)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}\leq\frac{K_{2}M^{\alpha}M_{1}^{\alpha}|\Omega|^{\alpha}}{\left(1-C_{SOR}(\Omega,N,p^{\prime})\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}\right)}\left\|\left(a+c\right)(\cdot+\tau,\cdot)\partial_{s}\phi_{\tau}(\cdot+\tau)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}^{1-\alpha}$
To obtain bounds which are independent of $\tau$, let us take
$\texttt{C}:=\frac{K_{2}MM_{1}^{\alpha}|\Omega|^{\alpha}}{\left(1-C_{SOR}(\Omega,N,p^{\prime})\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}\right)}\quad\mbox{ and }\quad\beta_{n}:=\left\|a+c\right\|_{\mathrm{L}^{p}(\Omega_{n,n+1})}\,\mbox{ for }n\in\mathbb{N}\cup\{0\}.$
As $\tau\geq 0$ is arbitrary, we deduce from (23) that
$\beta_{n+1}\leq\texttt{C}\,\beta_{n}^{1-\alpha}\qquad\mbox{ for }n\in\mathbb{N}.$
Consider the set
$\Lambda:=\big{\{}n\in\mathbb{N}\quad\text{ such that }\quad\beta_{n}\leq\beta_{n+1}\big{\}}.$
Observe that $\beta_{n}\leq\ {\texttt{C}}^{\frac{1}{\alpha}}$ for all $n\in\Lambda$. Hence we deduce that
$\beta_{n}\leq\max\big{\{}\beta_{0},{\texttt{C}}^{\frac{1}{\alpha}}\big{\}}.$
Observe that the sum of concentrations $a+c$ satisfies the following differential equation:
$\partial_{t}(a+c)-\Delta(\mu(a+c))&=0\qquad\qquad\qquad\qquad\ \mbox{in}\ \Omega_{T},\\ \nabla_{x}(a+c)\cdot n&=0\qquad\qquad\qquad\qquad\ \mbox{on}\ [0,T]\times\partial\Omega,\\ (a+c)(0,x)&=a_{0}+c_{0}\in{\mathrm{L}^{p}(\Omega)}\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
where $\mu:\Omega_{T}\to\mathbb{R}$ defined as follows:
$\mu(t,x):=\left(\frac{d_{a}a+d_{c}c}{a+c}\right)(t,x)\qquad\mbox{ for }(t,x)\in\Omega_{T}.$
Observe that $\mu$ satisfies
$0<\min\{d_{a},d_{c}\}\leq\mu(t,x)\leq\max\{d_{a},d_{c}\}\qquad\mbox{ for all }(t,x)\in\Omega_{T}.$
As the diffusion coefficients $d_{a}$ and $d_{c}$ satisfy the closeness condition (10), employing the $p^{\textrm{th}}$ order integrability estimate [CDF14, Proposition 1.1, p.1186] (see Theorem 12 from the Appendix for the precise statement), we arrive at
$\beta_{0}=\|a+c\|_{\mathrm{L}^{p}(\Omega_{0,1})}\leq\left(1+\max\{d_{a},d_{c}\}\frac{|d_{c}-d_{a}|\,C^{PRC}_{\frac{d_{a}+d_{c}}{2},p^{\prime}}}{2-|d_{c}-d_{a}|\,C^{PRC}_{\frac{d_{a}+d_{c}}{2},p^{\prime}}}\right)\|a_{0}+c_{0}\|_{\mathrm{L}^{p}(\Omega)}$
Hence we deduce
$\beta_{n}\leq\max\left\{\left(1+\max\{d_{a},d_{c}\}\frac{|d_{c}-d_{a}|\,C^{PRC}_{\frac{d_{a}+d_{c}}{2},p^{\prime}}}{2-|d_{c}-d_{a}|\,C^{PRC}_{\frac{d_{a}+d_{c}}{2},p^{\prime}}}\right)\|a_{0}+c_{0}\|_{\mathrm{L}^{p}(\Omega)},\,{\texttt{C}}^{\frac{1}{\alpha}}\right\}.$
Hence there exists a constant $C_{0}$, independent of $\tau$, such that
$\|a\|_{\mathrm{L}^{p}(\Omega_{\tau,\tau+1})}+\|c\|_{\mathrm{L}^{p}(\Omega_{\tau,\tau+1})}\leq C_{0}\qquad\forall\tau>0.$
The following lemma derives the key integrability estimate for the concentration $b$ which will play an important role in our analysis.
Lemma 2.
Let $p>N\geq 4$ and let $p^{\prime}$ be its Hölder conjugate. Let $(a,b,c)$ be the solution to the degenrate system (2) and let the nonzero diffusion coefficients $d_{a},d_{c}$ satisfy the closeness condition (10). Then there exists a constant $K_{3}>0$, depending only on the initial data and the dimension N, such that
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{\frac{N}{2}}(\Omega)}\leq K_{3}\,(1+t)^{\frac{N-2}{N-1}}\qquad\forall t\geq 0.$
Proof.
Recall that $b$ solves the equation
$\partial_{t}b=c-ab.$
Non-negativity of $a$ and $b$ implies that
$\partial_{t}b\leq c.$
Integrating over $(0,t)$ yields
$b(t,x)\leq b_{0}(x)+\int_{0}^{t}c(s,x)\,{\rm d}s$
Raising it to the power $N$ and employing Jensen’s inequality, we obtain
$\left(b(t,x)\right)^{N}\leq 2^{N-1}\left(\left(b_{0}(x)\right)^{N}+\left(\int_{0}^{t}c(s,x)\,{\rm d}s\right)^{N}\right)\leq 2^{N-1}(1+t)^{N-1}\left(\left(b_{0}(x)\right)^{N}+\int_{0}^{t}\left(c(s,x)\right)^{N}\,{\rm d}s\right)$
Integrating the above inequality in the $x$ variable over $\Omega$ yields
(24) $\left\|b(t,\cdot)\right\|_{\mathrm{L}^{N}(\Omega)}^{N}\leq 2^{N-1}(1+t)^{N-1}\Big{(}\left\|b_{0}\right\|_{\mathrm{L}^{N}(\Omega)}^{N}+\left\|c\right\|_{\mathrm{L}^{N}(\Omega_{t})}^{N}\Big{)}$
According to Lemma 1, for any $p>N$ and for any $\tau\geq 0$, we have
$\left\|c\right\|_{\mathrm{L}^{p}(\Omega_{\tau,\tau+1})}\leq C_{0},$
where the constant $C_{0}>0$ depends only on the dimension $N$, the domain $\Omega$ and the initial data. Hölder inequality yields
$\left\|c\right\|_{\mathrm{L}^{N}(\Omega_{\tau,\tau+1})}\leq\left\|c\right\|_{\mathrm{L}^{p}(\Omega_{\tau,\tau+1})}\,\left|\Omega\right|^{\frac{p-N}{p}}$
Hence we deduce that for any $t>0$,
$\left\|c\right\|_{\mathrm{L}^{N}(\Omega_{t})}^{N}\leq C_{0}^{N}\,\left|\Omega\right|^{\frac{N(p-N)}{p}}\,(1+t).$
Using this in (24) yields
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{N}(\Omega)}^{N}\leq 2^{N-1}(1+t)^{N-1}\left(\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{N}\,\left|\Omega\right|+C_{0}^{N}\,\left|\Omega\right|^{\frac{N(p-N)}{p}}\,(1+t)\right)$
Define a constant $K_{3}>0$ as follows:
$K_{3}^{\frac{N(N-1)}{N-2}}:=2^{N-1}\left(M_{2}|\Omega|\right)^{\frac{N}{N-2}}\left(\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{N}\,\left|\Omega\right|+C_{0}^{N}\,\left|\Omega\right|^{\frac{N(p-N)}{p}}\right)$
Thus we arrive at
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{N}(\Omega)}\leq K_{3}^{\frac{N-1}{N-2}}\left(M_{2}|\Omega|\right)^{-\frac{1}{N-2}}(1+t)$
Note that
$\frac{2}{N}=\frac{\alpha}{N}+\frac{1-\alpha}{1}\qquad\mbox{ with }\,\alpha=\frac{N-2}{N-1}.$
Hence by interpolation, we have
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{\frac{N}{2}}(\Omega)}\leq\left\|b(t,\cdot)\right\|_{\mathrm{L}^{N}(\Omega)}^{\frac{N-2}{N-1}}\,\left\|b(t,\cdot)\right\|_{\mathrm{L}^{1}(\Omega)}^{\frac{1}{N-1}}$
This yields the following estimate:
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{\frac{N}{2}}(\Omega)}\leq K_{3}\,(1+t)^{\frac{N-2}{N-1}}.$
Unlike the case of dimension $N\geq 4$, we are able to obtain $\mathrm{L}^{\frac{3}{2}}(\Omega)$-norm estimates on $b(t,\cdot)$ without any closeness assumption on the diffusion coefficients $d_{a}$ and $d_{c}$ in the case of dimensions $N\leq 3$. Next, we prove such an estimate and more in dimension three followed by a similar result in dimensions one and two.
Proposition 3.
Let $(a,b,c)$ be the solution to the degenerate system (2) in dimension $N=3$. Then, there exist constants $\hat{K},K_{c}>0$ and $\mu_{c}\in\mathbb{N}$, independent of time, such that
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}$ $\leq\hat{K}(1+t)^{\frac{5}{6}}\qquad\mbox{ for }\,t\geq 0,$
$\left\|a\right\|_{\mathrm{L}^{\frac{7}{2}}(\Omega_{\tau,\tau+1})}$ $\leq K_{c}\qquad\mbox{ for }\,\tau\geq 0,$
$\left\|c\right\|_{\mathrm{L}^{\frac{7}{2}}(\Omega_{\tau,\tau+1})}$ $\leq K_{c}(1+t)^{\mu_{c}}\qquad\mbox{ for }\,\tau\geq 0,$
Proof.
Recall that the solution $(a,b,c)$ to the degenerate system (2) satisfies the following estimate:
$\int_{0}^{t}\int_{\Omega}\left(4d_{a}\left|\nabla\sqrt{a}\right|^{2}+4d_{c}\left|\nabla\sqrt{c}\right|^{2}+(ab-c)\ln\left(\frac{ab}{c}\right)\right)\,{\rm d}x\,{\rm d}s\leq E(a_{0},b_{0},c_{0})-E(a,b,c)\leq E(a_{0},b_{0},c_{0}),$
where $E$ is the entropy functional defined in (9) and $(a_{\infty},b_{\infty},c_{\infty})$ is the homogeneous equilibrium state defined by (6)-(8). The above estimate along with the mass conservation properties (4)-(5) results in
(25) $\int_{\tau}^{\tau+2}\left\|\sqrt{a}(t,\cdot)\right\|_{\mathrm{W}^{1,2}(\Omega)}^{2}\,{\rm d}t\leq K_{4}\quad\mbox{ and }\quad\int_{\tau}^{\tau+2}\left\|\sqrt{c}(t,\cdot)\right\|_{\mathrm{W}^{1,2}(\Omega)}^{2}\,{\rm d}t\leq K_{4}\qquad\mbox{ for any }\tau\geq 0,$
where $K_{4}>0$ is the following constant
$K_{4}:=\frac{E(a_{0},b_{0},c_{0})}{4d_{a}}+\frac{E(a_{0},b_{0},c_{0})}{4d_{c}}+2M_{1}\left|\Omega\right|+2M_{2}\left|\Omega\right|.$
Gagliardo-Nirenberg inequality says that there exist constants $K_{5}=K_{5}(\Omega,N)$ and $K_{6}=K_{6}(\Omega,N,r,q)$ where $N$ is the dimension and $1\leq p,q,r\leq+\infty$ such that for nonnegative integers $j<m$,
$\left\|D^{j}u\right\|_{\mathrm{L}^{p}(\Omega)}\leq K_{5}\left\|D^{m}u\right\|_{\mathrm{L}^{r}(\Omega)}^{\alpha}\left\|u\right\|_{\mathrm{L}^{q}(\Omega)}^{1-\alpha}+K_{6}\left\|u\right\|_{\mathrm{L}^{q}(\Omega)}\quad\mbox{ for any }u\in\mathrm{L}^{q}(\Omega)\mbox{ satisfying }D^{m}u\in\mathrm{L}^{r}(\Omega).$
Furthermore, the various parameters in the above inequality satisfy
$\frac{1}{p}=\frac{j}{N}+\big{(}\frac{1}{r}-\frac{m}{N}\big{)}\alpha\ +\frac{1-\alpha}{q}\qquad\mbox{ and }\qquad\frac{j}{N}\leq\alpha\leq 1.$
Applying the Gagliardo-Nirenberg inequality for $\sqrt{a}$ by taking $r=2,m=1,j=0,N=3,q=2$ and $\alpha=\frac{3}{5}$ yields
$\left\|\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{\frac{10}{3}}(\Omega)}\leq K_{5}\left\|\nabla\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}^{\frac{3}{5}}\,\left\|\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}^{\frac{2}{5}}+K_{6}\left\|\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}$
Raising it to the power $\frac{10}{3}$ yields
$\left\|\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{\frac{10}{3}}(\Omega)}^{\frac{10}{3}}$ $\leq 2^{\frac{7}{3}}K_{5}^{\frac{10}{3}}\left\|\nabla\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}^{2}\,\left\|\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}^{\frac{4}{3}}+2^{\frac{7}{3}}K_{6}^{\frac{10}{3}}\left\|\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}^{\frac{10}{3}}$
$=2^{\frac{7}{3}}K_{5}^{\frac{10}{3}}\left\|\nabla\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}^{2}\,M_{1}^{\frac{4}{3}}\,\left|\Omega\right|^{\frac{4}{3}}+2^{\frac{7}{3}}K_{6}^{\frac{10}{3}}\,M_{1}^{\frac{10}{3}}\,\left|\Omega\right|^{\frac{10}{3}}$
Take an arbitrary $\tau\geq 0$ and integrate the above inequality from $\tau$ to $\tau+2$ in the $t$ variable to get
$\int_{\tau}^{\tau+2}\left\|\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{\frac{10}{3}}(\Omega)}^{\frac{10}{3}}\,{\rm d}t\leq 2^{\frac{7}{3}}K_{5}^{\frac{10}{3}}\,K_{4}\,M_{1}^{\frac{4}{3}}\,\left|\Omega\right|^{\frac{4}{3}}+2^{\frac{10}{3}}K_{6}^{\frac{10}{3}}\,M_{1}^{\frac{10}{3}}\,\left|\Omega\right|^{\frac{10}{3}},$
where we have used the apriori bound (25). Similarly, we can obtain the following estimate for $\sqrt{c}$:
$\int_{\tau}^{\tau+2}\left\|\sqrt{c}(t,\cdot)\right\|_{\mathrm{L}^{\frac{10}{3}}(\Omega)}^{\frac{10}{3}}\,{\rm d}t\leq 2^{\frac{7}{3}}K_{5}^{\frac{10}{3}}\,K_{4}\,M_{1}^{\frac{4}{3}}\,\left|\Omega\right|^{\frac{4}{3}}+2^{\frac{10}{3}}K_{6}^{\frac{10}{3}}\,M_{1}^{\frac{10}{3}}\,\left|\Omega\right|^{\frac{10}{3}}.$
Define a constant $K_{7}>0$ as follows:
$K_{7}:=\left(2^{\frac{7}{3}}K_{5}^{\frac{10}{3}}\,K_{4}\,M_{1}^{\frac{4}{3}}\,\left|\Omega\right|^{\frac{4}{3}}+2^{\frac{10}{3}}K_{6}^{\frac{10}{3}}\,M_{1}^{\frac{10}{3}}\,\left|\Omega\right|^{\frac{10}{3}}\right)^{\frac{3}{5}}.$
We have thus arrived at the following integrability estimates:
(26) $\left\|a\right\|_{\mathrm{L}^{\frac{5}{3}}(\Omega_{\tau,\tau+2})}\leq K_{7}\quad\mbox{ and }\quad\left\|c\right\|_{\mathrm{L}^{\frac{5}{3}}(\Omega_{\tau,\tau+2})}\leq K_{7}\quad\mbox{ with }\tau\geq 0.$
Recall that non-negativity of $a,b$ implies $\partial_{t}b\leq c$ which in turn implies that
$b(t,x)\leq b_{0}(x)+\int_{0}^{t}c(s,x)\,{\rm d}s.$
Raising the above inequality to the power $\frac{5}{3}$ yields
$\left(b(t,x)\right)^{\frac{5}{3}}\leq 2^{\frac{2}{3}}(1+t)^{\frac{2}{3}}\left(\left(b_{0}(x)\right)^{\frac{5}{3}}+\int_{0}^{t}\left(c(s,x)\right)^{\frac{5}{3}}\,{\rm d}s\right).$
Integrating the above inequality over $\Omega$ yields
$\int_{\Omega}\left(b(t,x)\right)^{\frac{5}{3}}\,{\rm d}x$ $\leq 2^{\frac{2}{3}}(1+t)^{\frac{2}{3}}\left(\int_{\Omega}\left(b_{0}(x)\right)^{\frac{5}{3}}\,{\rm d}x+\int_{0}^{t}\int_{\Omega}\left(c(s,x)\right)^{\frac{5}{3}}\,{\rm d}x\,{\rm d}s\right)$
$\leq 2^{\frac{2}{3}}(1+t)^{\frac{2}{3}}\left(\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{\frac{5}{3}}\,\left|\Omega\right|+\sum_{\tau=0}^{1+\lfloor t\rfloor}\left\|c\right\|_{\mathrm{L}^{\frac{5}{3}}(\Omega_{\tau,\tau+1})}^{\frac{5}{3}}\right)$
$\leq 2^{\frac{2}{3}}(1+t)^{\frac{2}{3}}\left(\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{\frac{5}{3}}\,\left|\Omega\right|+(1+t)K_{7}^{\frac{5}{3}}\right)$
$\leq 2^{\frac{2}{3}}(1+t)^{\frac{5}{3}}\left(\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{\frac{5}{3}}\,\left|\Omega\right|+K_{7}^{\frac{5}{3}}\right)$
Define a constant $K_{8}>0$ as follows:
$K_{8}:=\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{\frac{5}{3}}\,\left|\Omega\right|+K_{7}^{\frac{5}{3}}.$
We have thus arrived at the following integrability estimate:
(27) $\left\|b(t,\cdot)\right\|_{\mathrm{L}^{\frac{5}{3}}(\Omega)}^{\frac{5}{3}}\leq K_{8}(1+t)^{\frac{5}{3}}\qquad\mbox{ for }t\geq 0.$
Note that
$\frac{2}{3}=\frac{3\alpha}{5}+\frac{1-\alpha}{1}\qquad\mbox{ with }\,\alpha=\frac{5}{6}.$
Hence by interpolation, we have
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}$ $\leq\left\|b(t,\cdot)\right\|_{\mathrm{L}^{\frac{5}{3}}(\Omega)}^{\frac{5}{6}}\,\left\|b(t,\cdot)\right\|_{\mathrm{L}^{1}(\Omega)}^{\frac{1}{6}}\leq K_{8}^{\frac{1}{2}}\,M_{2}^{\frac{1}{6}}\left|\Omega\right|^{\frac{1}{6}}(1+t)^{\frac{5}{6}},$
thanks to the estimate (27) and the mass conservation property (5).
As $a$ is a positive subsolution of the parabolic equation $\partial_{t}a-d_{a}\Delta a\leq c$, integrability estimation (11) yields
$\|a\|_{L^{4.5}(\Omega_{0,2})}\leq C_{IE}(\Omega,d_{a})\|c\|_{L^{\frac{5}{3}}(\Omega_{0,2})}\leq C_{IE}(\Omega,d_{a})K_{7}.$
Consider $0\leq\Theta\in C_{c}^{\infty}(\Omega_{0,2})$(space of all compactly supported smooth function) satisfies
$-[\partial_{t}\phi+d_{c}\Delta\phi]=$ $\Theta\qquad\ \ \ \Omega_{0,2}$
$\nabla\phi.n=$ $0\qquad(0,2)\times\partial\Omega$
ϕ(2)=italic-$\phi(2)=$ $0\qquad\ \ \ \ \Omega$
We have $\phi\geq 0$, and the following estimates for a constant $C_{q,d_{c}}>0$( $q\in(1,\infty)$, arbitrary)[Pie10][Ama85]
$\|\phi_{t}\|_{L^{q}(\Omega_{0,2})}+\|\Delta\phi\|_{L^{q}(\Omega_{0,2})}+\sup\limits_{s\in[0,2]}\|\phi(s)\|_{L^{q}(\Omega)}+\|\phi\|_{L^{q}((0,2)\times\partial\Omega)}\leq C_{q,d_{c}}\|\Theta\|_{L^{q}(\Omega_{0,2})}.$
We derive the following estimate of $c$, through integration by parts:
$\int_{\Omega_{0,2}}c\Theta=$ $\int_{\Omega}c_{0}\phi(0)-\int_{\Omega_{0,2}}(\partial_{t}-d_{c}\Delta)c\phi$
$=$ $\int_{\Omega}c_{0}\phi(0)+\int_{\Omega_{0,2}}(\partial_{t}-d_{a}\Delta)a\phi=\int_{\Omega}(c_{0}+a_{0})\phi(0)+\int_{\Omega_{0,2}}a\partial_{t}\phi+\int_{\Omega_{0,2}}ad_{a}\Delta\phi$
$\leq$ $\|a_{0}+c_{0}\|_{L^{p}(\Omega)}\|\phi(0)\|_{L^{q}(\Omega)}+\|a\|_{L^{p}(\Omega_{0,2})}\|\partial_{t}\phi\|_{L^{q}(\Omega_{0,2})}+d_{a}\|a\|_{L^{p}(\Omega_{0,2})}\|\Delta\phi\|_{L^{q}(\Omega_{0,2})}.$
Choose $q=\frac{4.5}{4.5-1}$,i.e., Hölder conjugate of $4.5$,duality estimate yields
$\|c\|_{L^{4.5}(\Omega_{0,2})}\leq\|a_{0}+c_{0}\|_{L^{4.5}(\Omega)}+2C_{q,d_{c}}\|a\|_{L^{4.5}(\Omega_{0,2})}.$
Choose $K_{0,1}=\max\left\{(C_{IE}+2C_{q,d_{c}})(\Omega,d_{a})K_{7}+\|a_{0}+c_{0}\|_{L^{4.5}(\Omega)}\right\}$, we have the following estimate
$\|a\|_{L^{4.5}(\Omega_{0,1})}+\|c\|_{L^{4.5}(\Omega_{0,1})}\leq K_{0,1}.$
Let $\phi:[0,\infty)\rightarrow[0,1]$ be a smooth function such that $\phi(0)=0$ and
$\phi(x)$ $=1\qquad\mbox{ for }x\in[1,\infty)$
$0\leq\phi^{\prime}(x)$ $\leq M\qquad\mbox{ for }x\in[0,\infty)$
for some constant $M>0$. For an arbitrary $\tau>0$, consider $\phi_{\tau}:[\tau,\infty)\to[0,1]$ defined as $\phi_{\tau}(s):=\phi(s-\tau)$ for $s\in[\tau,\infty)$. Then, the product $\phi_{\tau}(t)a(t,x)$ satisfies
$\partial_{t}\left(\phi_{\tau}a\right)-d_{a}\Delta\left(\phi_{\tau}a\right)&=a\partial_{t}\phi_{\tau}+\phi_{\tau}(c-ab)\qquad\mbox{ in}\Omega_{\tau,\tau+2}\\ \phi_{\tau}\nabla a\cdot n&=0\qquad\qquad\qquad\qquad\qquad\ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}(\tau)a(\tau,x)&=0\qquad\qquad\qquad\qquad\qquad\ \ \ \mbox{ in }\Omega\end{aligned}\right.$
Let $\zeta(t,x)$ be the solution to the following initial boundary value problem:
[…]
By employing the maximum principle for the heat equation and exploiting the positivity of $a,b$, we deduce
(28) $\phi_{\tau}(t,x)a(t,x)\leq\zeta(t.x)\qquad\mbox{ for }(t,x)\in\Omega_{\tau,\tau+2}.$
Employing the integrability estimate for $\zeta$ available from [CDF14, Lemma 3.3] (see Theorem 11 from the Appendix for the precise statement), we deduce that
$\left\|\zeta\right\|_{\mathrm{L}^{s}(\Omega_{\tau,\tau+2})}\leq C_{IE}(\Omega,d_{a},s)\left\|a\partial_{t}\phi_{\tau}+\phi_{\tau}c\right\|_{\mathrm{L}^{\frac{5}{3}}(\Omega_{\tau,\tau+2})}\leq 2C_{IE}(\Omega,d_{a},s)(1+M)K_{7}\qquad\mbox{ for any }s<5,$
where the second inequality is a consequence of the properties of smooth function $\phi$ and the integrability estimates on $a$ and $c$ from (26). Thanks to the pointwise bound (28), we in particular (taking $s=\frac{9}{2}$) have
$\left\|a\right\|_{\mathrm{L}^{\frac{9}{2}}(\Omega_{\tau+1,\tau+2})}\leq\left\|\zeta\right\|_{\mathrm{L}^{\frac{9}{2}}(\Omega_{\tau,\tau+2})}\leq 2C_{IE}(\Omega,d_{a})(1+M)K_{7}\qquad\mbox{ for }\tau\geq 0.$
Hence we deduce for $\tau\geq 0$,
(29) $\left\|a\right\|_{\mathrm{L}^{\frac{9}{2}}(\Omega_{\tau,\tau+1})}\leq\left\|a\right\|_{\mathrm{L}^{\frac{9}{2}}(\Omega_{0,1})}+2C_{IE}(\Omega,d_{a})(1+M)K_{7}\leq K_{0,1}+2C_{IE}(\Omega,d_{a})(1+M)K_{7}=:K_{9}.$
Furthermore, integrating the estimate (27) in the time variable from $\tau$ to $\tau+1$ helps us obtain
$\left\|b\right\|_{\mathrm{L}^{\frac{5}{3}}(\Omega_{\tau,\tau+1})}\leq 2K_{8}^{\frac{3}{5}}(1+\tau)=:K_{10}(1+\tau)\qquad\mbox{ for }\tau\geq 0.$
Observe that an application of the Hölder inequality yields
$\left\|ab\right\|_{\mathrm{L}^{\frac{45}{37}}(\Omega_{\tau,\tau+1})}\leq\left\|a\right\|_{\mathrm{L}^{\frac{9}{2}}(\Omega_{\tau,\tau+1})}^{\frac{2}{9}}\,\left\|b\right\|_{\mathrm{L}^{\frac{5}{3}}(\Omega_{\tau,\tau+1})}^{\frac{3}{5}}\leq K_{9}^{\frac{2}{9}}K_{10}^{\frac{3}{5}}(1+\tau)^{\frac{3}{5}}$
Using the smooth function $\phi$ defined earlier, remark that the product $\phi_{\tau}(t)c(t,x)$ satisfies
$\partial_{t}\left(\phi_{\tau}c\right)-d_{c}\Delta\left(\phi_{\tau}c\right)=&c\partial_{t}\phi_{\tau}+\phi_{\tau}(ab-c)\qquad\mbox{ in}\Omega_{\tau,\tau+2}\\ \phi_{\tau}\nabla c\cdot n&=0\qquad\qquad\qquad\qquad\ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}(\tau)c(\tau,x)&=0\qquad\qquad\qquad\qquad\ \ \ \mbox{ in }\Omega\end{aligned}\right.$
where $\phi_{\tau}(s):=\phi(s-\tau)$ for $s\in[\tau,\infty)$ with $\tau>0$. Thanks to the positivity of $c$, employing the maximum principle for the heat equation and the integrability estimate from [CDF14, Lemma 3.3] (see Theorem 11 from the Appendix for the precise statement), we arrive at
$\left\|\phi_{\tau}c\right\|_{\mathrm{L}^{s}(\Omega_{\tau,\tau+2})}$ $\leq C_{IE}(\Omega,d_{c},s)\left\|c\partial_{t}\phi_{\tau}+\phi_{\tau}ab\right\|_{\mathrm{L}^{\frac{45}{37}}(\Omega_{\tau,\tau+2})}$
$\leq C_{IE}\left\{M\left\|c\right\|_{\mathrm{L}^{\frac{45}{37}}(\Omega_{\tau,\tau+2})}+\left\|ab\right\|_{\mathrm{L}^{\frac{45}{37}}(\Omega_{\tau,\tau+2})}\right\}$
$\leq C_{IE}\left\{M\left\|c\right\|^{\frac{5}{9}}_{\mathrm{L}^{1}(\Omega_{\tau,\tau+2})}\left\|c\right\|^{\frac{4}{9}}_{\mathrm{L}^{\frac{5}{3}}(\Omega_{\tau,\tau+2})}+\left\|ab\right\|_{\mathrm{L}^{\frac{45}{37}}(\Omega_{\tau,\tau+2})}\right\}$
$\leq C_{IE}\left\{MM_{1}^{\frac{5}{9}}K_{7}^{\frac{4}{9}}+2K_{9}^{\frac{2}{9}}K_{10}^{\frac{3}{5}}(1+\tau)^{\frac{3}{5}}\right\}$
for any $s<\frac{(5)\frac{45}{37}}{(5-\frac{90}{37})}$. Taking $s=\frac{9}{4}$ leads to
$\left\|\phi_{\tau}c\right\|_{\mathrm{L}^{\frac{9}{4}}(\Omega_{\tau,\tau+2})}\leq K_{11}(1+\tau)^{\frac{3}{5}}$
where $K_{11}:=C_{IE}(\Omega,d_{c})\max\left\{MM_{1}^{\frac{5}{9}}K_{7}^{\frac{4}{9}},2K_{9}^{\frac{2}{9}}K_{10}^{\frac{3}{5}}\right\}$. Hence we deduce
$\left\|c\right\|_{\mathrm{L}^{\frac{9}{4}}(\Omega_{\tau,\tau+1})}\leq\left\|c\right\|_{\mathrm{L}^{\frac{9}{4}}(\Omega_{0,1})}+K_{11}(1+\tau)^{\frac{3}{5}}\leq K_{0,1}\left|\Omega\right|^{\frac{1}{2}}+K_{11}(1+\tau)^{\frac{3}{5}}\qquad\mbox{ for }\tau\geq 0.$
Again, exploiting the relation $\partial_{t}b\leq c$, we arrive at
$\left\|b(t,\cdot)\right\|^{\frac{9}{4}}_{\mathrm{L}^{\frac{9}{4}}(\Omega)}$ $\leq 2^{\frac{5}{4}}(1+t)^{\frac{5}{4}}\left(\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{\frac{9}{4}}\,\left|\Omega\right|+(1+t)^{\frac{47}{20}}K_{12}^{\frac{9}{4}}\right)\leq K_{13}(1+t)^{\frac{18}{5}}$
thus deducing
$\left\|b\right\|_{\mathrm{L}^{\frac{9}{4}}(\Omega_{\tau,\tau+1})}\leq\left(\frac{5K_{13}}{23}\right)^{\frac{4}{9}}(1+\tau)^{\frac{92}{45}}\qquad\mbox{ for }\tau\geq 0.$
Hölder inequality results in
$\left\|ab\right\|_{\mathrm{L}^{\frac{3}{2}}(\Omega_{\tau,\tau+1})}\leq\left\|a\right\|_{\mathrm{L}^{\frac{9}{2}}(\Omega_{\tau,\tau+1})}\left\|b\right\|_{\mathrm{L}^{\frac{9}{4}}(\Omega_{\tau,\tau+1})}\leq K_{9}\left(\frac{5K_{13}}{23}\right)^{\frac{4}{9}}(1+\tau)^{\frac{92}{45}},$
where we have used the bounds obtained earlier. As $\frac{7}{2}<\frac{(5)\frac{3}{2}}{5-3}$, employing the maximum principle for the heat equation and the integrability estimate from [CDF14, Lemma 3.3] (see Theorem 11 from the Appendix for the precise statement) as before, we get
$\left\|c\right\|_{\mathrm{L}^{\frac{7}{2}}(\Omega_{\tau,\tau+1})}\leq K_{14}(1+\tau)^{\frac{92}{45}}\qquad\mbox{ for }\tau\geq 0.$
In our above proof Gagliardo-Nirenberg inequality played an important role. We will use the same technique to get this   $L^{\frac{3}{2}}(\Omega)$ integral estimation of  $b$  for N=1,2 case.
Lemma 3.
Let $a,.b,c$ be the solution of the degenerate system-(2) in dimension $N=1,2$, then there exists constants $\hat{K},K_{c}>0$ and some $\mu_{c}\in\mathbb{N}$, independent of time, such that:
[…]
Proof.
Similar as in the proof of proposition-3, we are going to use Gagliardo-Nirenberg inequality which says there exists a constant depending on the domain $C_{2}=C_{2}(\Omega,N)$ N be the dimension , for $1\leq p,q,r\leq+\infty$
$\|D^{j}\sqrt{a}\|_{L^{p}(\Omega)}\leq C_{2}\|D^{m}\sqrt{a}\|_{L^{r}(\Omega)}^{\alpha}\|\sqrt{a}\|_{L^{q}(\Omega)}^{1-\alpha}+\tilde{C_{2}}(\Omega,N,r,q)\|\sqrt{a}\|_{L^{q}(\Omega)}$
where,   $\frac{1}{p}=\frac{j}{N}+\big{(}\frac{1}{r}-\frac{m}{N}\big{)}\alpha\ +\frac{1-\alpha}{q}$   and   $\frac{j}{N}\leq\alpha\leq 1$
For N=2 take  $r=2,m=1,j=0,q=2$ and $\alpha=\frac{1}{2}$ we get
$\|\sqrt{a}\|_{L^{4}(\Omega)}^{4}\leq 2^{3}C_{2}^{4}\|\nabla\sqrt{a}\|_{L^{2}(\Omega)}^{2}\|\sqrt{a}\|_{L^{2}(\Omega)}^{2}+2^{3}\tilde{C_{2}}^{4}\|\sqrt{a}\|_{L^{2}(\Omega)}^{4}$
We have just like in theorem 3   $\forall\tau\geq 0$
$\int_{\tau}^{\tau+2}\|\sqrt{a}\|_{W^{1,2}(\Omega)}^{2}\leq K_{1};\ \ \int_{\tau}^{\tau+2}\|\sqrt{c}\|_{W^{1,2}(\Omega)}^{2}\leq K_{1}$
Where $K_{1}=\frac{2E(a_{0},b_{0},c_{0})}{d_{a}}+\frac{2E(a_{0},b_{0},c_{0})}{d_{c}}+2M_{1}\Omega+2M_{2}\Omega$. Integrating from  $\tau$  to  $\tau+2$
$\int_{\tau}^{\tau+2}\|\sqrt{a}\|_{L^{4}(\Omega)}^{4}\leq 2^{4}C_{2}^{4}M_{1}^{2}|\Omega|^{2}K_{1}+2^{4}\tilde{C_{2}}^{4}M_{1}^{4};$
$\int_{\tau}^{\tau+2}\|\sqrt{c}\|_{L^{4}(\Omega)}^{4}\leq 2^{4}C_{2}^{4}M_{1}^{2}|\Omega|^{2}K_{1}+2^{4}\tilde{C_{2}}^{4}M_{1}^{4}$
Take  $K_{2}=(2^{4}C_{2}^{4}M_{1}^{2}|\Omega|^{2}K_{1}+2^{4}\tilde{C_{2}}^{4}M_{1}^{4})^{\frac{1}{2}}$, we have, $\forall\tau\geq 0$
$\|a\|_{L^{2}(\Omega_{\tau,\tau+2})}\leq K_{2},\qquad\|c\|_{L^{2}(\Omega_{\tau,\tau+2})}\leq K_{2}$
Applying Hölder inequality we get for all  $\tau\geq 0$,
(30) $\|a\|_{L^{\frac{5}{3}}(\Omega_{\tau,\tau+2})}\leq 2K_{2}|\Omega|^{\frac{1}{6}},\qquad\|c\|_{L^{\frac{5}{3}}(\Omega_{\tau,\tau+2})}\leq 2K_{2}|\Omega|^{\frac{1}{6}}$
Rest of the proof is just like the proof of proposition 3. We will get  $\hat{K}>0$ such that for  $t\geq 0$
$\|b\|_{L^{\frac{3}{2}}(\Omega)}\leq\hat{K}(1+t)^{\frac{5}{6}}$
Now for N=1 we take, $r=2,m=1,j=0,q=2$ and $\alpha=\frac{1}{3}$ and like in the previous N=2 case a similar approach gives us $\exists\ \hat{K}>0,K_{c}>0$ and $\mu_{c}\in\mathbb{N}$, such that for $\tau\geq 0$
$\|b\|_{L^{\frac{3}{2}}(\Omega)}$ $\leq\hat{K}(1+\tau)^{\frac{5}{6}};\qquad\ \tau\geq 0$
$\|a\|_{L^{3.5}(\Omega_{\tau,\tau+1})}$ $\leq K_{c}(1+\tau)^{\mu_{c}};\qquad\tau\geq 0$
$\|c\|_{L^{3.5}(\Omega_{\tau,\tau+1})}$ $\leq K_{c}(1+\tau)^{\mu_{c}};\qquad\tau\geq 0$
Now above proposition-1 and proposition-3 will help us to relate entropy dissipation with the missing term  $\delta_{B}$. Furthermore the integral estimates of $c$ will provide us polynomial growth for the solutions too. We begin by the polynomial growth in the next theorem.
Lemma 4.
Let $a,b,c$ be the solution of degenerate system-(2). Furthermore let for $p>\frac{N+2}{2}$, there exists a constant $K_{c}>0$, and some $\mu_{c}\in\mathbb{N}$, independent of time, such that:
$\|a\|_{L^{p}(\Omega_{\tau,\tau+1})}+\|c\|_{L^{p}(\Omega_{\tau,\tau+1})}\leq K_{c}(1+\tau)^{\mu_{c}},\quad\forall\tau\geq 0$
Then there exists constant $K_{\infty}$ and $\mu\in\mathbb{N}$, such that:
[…]
Proof.
Cnsidering $\phi:[0,\infty)\rightarrow[0,1]$ be a smooth function such that $\phi(0)=0$; $\phi\big{|}_{[1,\infty)}=1\text{\ and \ }\phi^{\prime}\in[0,M_{\phi}]$.
Let’s denote $\phi_{\tau}(s)=\phi(s-\tau)$, then:
$\partial_{t}\phi_{\tau}a-d_{a}\Delta\phi_{\tau}a=&\phi^{\prime}_{\tau}a+\phi_{\tau}(c-ab)\qquad\mbox{ in }\Omega_{\tau,\tau+2}\\ \nabla\phi_{\tau}a.\gamma=&0\qquad\qquad\ \ \qquad\qquad\ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}a(\tau,x)=&0\qquad\qquad\qquad\ \qquad\ \ \ \mbox{ in }\Omega\end{aligned}\right.$
Then from integrability estimation(appendix-11) we have:
$\|\phi_{\tau}a\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\leq$ $C_{IE}(\Omega,N,p)\|\phi^{\prime}_{\tau}a+\phi_{\tau}c\|_{L^{p}(\Omega_{\tau,\tau+1})}$
(31) $\|a\|_{L^{\infty}(\Omega_{\tau+1,\tau+2})}\leq$ $C_{IE}(\Omega,N,p)(1+M_{\phi})K_{c}(1+\tau)^{\mu_{c}}$
Furthermore from the relation $partial_{t}b\leq c$ using jensen inequality we get for some $\mu_{1}\in\mathbb{N}$
$\|b\|_{L^{p}(\Omega)}\leq K_{1}(1+t)^{\mu_{1}}$
Integrating with respect to time variable from $\tau$ to $\tau+1$, we will get some $\mu_{2}\in\mathbb{N}$
$\|b\|_{L^{p}(\Omega_{\tau,\tau+1})}\leq K_{2}(1+\tau)^{\mu_{2}}$
This along with relation-(31) implies for some $\mu_{3}\mathbb{N}$
$\|ab\|_{L^{p}(\Omega_{\tau,\tau+1})}\leq K_{3}(1+\tau)^{\mu_{3}};\qquad p>\frac{N+2}{2}$
Considering the same smooth cutoff function $\phi$
$\partial_{t}\phi_{\tau}c-d_{c}\Delta\phi_{\tau}c=&\phi^{\prime}_{\tau}c+\phi_{\tau}(ab-c)\qquad\mbox{ in }\Omega_{\tau,\tau+2}\\ \nabla\phi_{\tau}c.\gamma=&0\qquad\qquad\ \ \qquad\qquad\ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}c(\tau,x)=&0\qquad\qquad\qquad\ \qquad\ \ \ \mbox{ in }\Omega\end{aligned}\right.$
Then from integrability estimation(appendix-11) we have:
$\|\phi_{\tau}c\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\leq$ $C_{IE}(\Omega,N,p)\|\phi^{\prime}_{\tau}c+\phi_{\tau}ab\|_{L^{p}(\Omega_{\tau,\tau+1})}$
(32) $\|c\|_{L^{\infty}(\Omega_{\tau+1,\tau+2})}\leq$ $C_{IE}(\Omega,N,p)(1+M_{\phi})(K_{c}+K_{3})(1+\tau)^{\mu_{c}+\mu_{3}}$
Now again from the relation $\partial_{t}b\leq c$ using jensen inequality we get for some $\mu_{4}\in\mathbb{N}$
$\|b\|_{L^{\infty}(\Omega_{t})}\leq K_{4}(1+t)^{\mu_{4}}$
This along with supremum estimates of $a$((31)) and $c$((32)) gives us the required result, that there exists constant $K_{\infty}$ and $\mu\in\mathbb{N}$, such that:
[…]
Note for $N\geq 4$, if $d_{a},d_{c}$ satisfying closeness condition for $p>N\geq\frac{N+2}{2}$, the assumption in the theorem-4 is automatically satisfied(proposition-1), however the assumption holds true also for $N=1,2,3$ without any closeness condition. ∎
Next we show the integral estimate of $b$, as in lemma-2 and proposition-3, helps us to relate the missing term $\delta_{B}$ with the entropy dissipation functional. The relation is described in the following proposition:
Proposition 4.
Let $N\geq 4$ and let $(a,b,c)$ be the solution to the degenrate system (2). Let the nonzero diffusion coefficients $d_{a},d_{c}$ satisfy the closeness condition (10). Then the entropy dissipation $D(a,b,c)$ satisfies
(33) $D(a,b,c)\geq\mathcal{K}\,(1+t)^{-\frac{N-2}{N-1}}\left(\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\right)\qquad\mbox{ for }t\geq 0,$
where the positive constant $\mathcal{K}$ depends only on the dimension N, the domain $\Omega$, the constants $M_{1}$ and $M_{2}$ in the mass conservation properties (4)-(5) and the nonzero diffusion coefficients $d_{a}$ and $d_{c}$.
Proof.
We rewrite the entropy dissipation functional as
(34) $D(a,b,c)=4d_{a}\int_{\Omega}\left|\nabla A\right|^{2}\,{\rm d}x+4d_{c}\int_{\Omega}\left|\nabla C\right|^{2}\,{\rm d}x+\int_{\Omega}(ab-c)\ln\left(\frac{ab}{c}\right)\,{\rm d}x.$
We recall an algebraic identity which says that for all $p,q\geq 0$, there holds $(p-q)\left(\ln p-\ln q\right)\geq 4\left(\sqrt{p}-\sqrt{q}\right)^{2}$. Using this algebraic identity in the last term and employing the Poincaré-Wirtinger inequality (see Theorem 7 from the Appendix for the precise statement) for the first two terms of the above dissipation functional yields
(35) $D(a,b,c)\geq\frac{4d_{a}}{P(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}+\frac{4d_{c}}{P(\Omega)}\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}+4\left\|AB-C\right\|^{2}_{\mathrm{L}^{2}(\Omega)},$
where $P(\Omega)=C\left(\Omega,\frac{2N}{N-2}\right)$ is the Poincaré constant. From the above inequality, it follows that
(36) $D(a,b,c)\geq\frac{4d_{a}}{P(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}+\frac{4d_{c}}{P(\Omega)}\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}+\eta\left\|AB-C\right\|^{2}_{\mathrm{L}^{2}(\Omega)},$
for any $0\leq\eta\leq 4$. As a consequence of the Hölder inequality, the inequality (35) leads to
$D(a,b,c)\geq\frac{4d_{a}\left|\Omega\right|^{-\frac{2}{N}}}{P(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{4d_{c}\left|\Omega\right|^{-\frac{2}{N}}}{P(\Omega)}\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+4\left\|AB-C\right\|^{2}_{\mathrm{L}^{2}(\Omega)}.$
It is apparent from the above lower bound that a term involving $\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$ is missing. To arrive at a lower bound involving this missing term, we fix an arbitrary constant $\varepsilon>0$ (to be chosen later) and distinguish two cases: a case corresponding to
(37) $\max\left\{\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)},\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\right\}\leq\varepsilon$
and another case corresponding to
(38) $\max\left\{\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)},\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\right\}>\varepsilon.$
We first treat the case corresponding to (37). Observe that
$\left\|AB-C\right\|^{2}_{\mathrm{L}^{2}(\Omega)}=\left\|\left(\delta_{A}+\overline{A}\right)B-\left(\delta_{C}+\overline{C}\right)\right\|^{2}_{\mathrm{L}^{2}(\Omega)}=\left\|\left(\,\overline{A}B-\overline{C}\right)+\left(B\delta_{A}-\delta_{C}\right)\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$
We recall an algebraic identity which says that for all $p,q\in\mathbb{R}$, there holds $(p-q)^{2}\geq\frac{p^{2}}{2}-q^{2}$. Using this algebraic identity in the above equality, we obtain
$\left\|AB-C\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$ $\geq\frac{1}{2}\left\|\overline{A}B-\overline{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}-\left\|B\delta_{A}-\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$
$\geq\frac{1}{2}\left\|\overline{A}B-\overline{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}-2\left\|B\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}-2\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)},$
thanks to the algebraic identity $(p-q)^{2}\leq 2p^{2}+2q^{2}$ for $p,q\in\mathbb{R}$. Employing the Hölder inequality in the second and third terms of the lower bound in the above inequality results in
$\left\|AB-C\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\geq\frac{1}{2}\left\|\overline{A}B-\overline{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}-2\left\|b\right\|_{\mathrm{L}^{\frac{N}{2}}(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}-2\left|\Omega\right|^{\frac{2}{N}}\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}.$
Next, using the key integrability estimate on $\left\|b(t,\cdot)\right\|_{\mathrm{L}^{\frac{N}{2}}(\Omega)}$ from Lemma 2 yields
(39) $\left\|AB-C\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\geq\frac{1}{2}\left\|\overline{A}B-\overline{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}-2K_{3}\,(1+t)^{\frac{N-2}{N-1}}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}-2\left|\Omega\right|^{\frac{2}{N}}\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}.$
In order to relate the above lower bound to $\left\|\delta_{B}\right\|_{\mathrm{L}^{2}(\Omega)}$, we further analyse the following term:
$\left\|\overline{A}B-\overline{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}.$
Note that if $\overline{A}\geq\sqrt{\varepsilon}$, then
(40) $\left\|\overline{A}B-\overline{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\geq\varepsilon\left\|B-\overline{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}.$
To see this, factorising $\overline{A}B$ as $\overline{C}(1+\mu(x))$, we get
$\left\|\overline{A}B-\overline{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}=\overline{C}^{2}\overline{\mu^{2}}\left|\Omega\right|\quad\mbox{ and }\quad\left\|B-\overline{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}=\frac{\overline{C}^{2}}{\overline{A}^{2}}\left\|\mu-\overline{\mu}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\leq\frac{\overline{C}^{2}}{\overline{A}^{2}}\overline{\mu^{2}}\left|\Omega\right|\leq\frac{\overline{C}^{2}\overline{\mu^{2}}\left|\Omega\right|}{\varepsilon}.$
On the other hand, let us consider the case when $\overline{A}<\sqrt{\varepsilon}$. Note that
$\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}=\left|\Omega\right|\left(\overline{A^{2}}-\overline{A}^{2}\right)\implies\overline{A^{2}}\leq\varepsilon\left(1+\frac{1}{\left|\Omega\right|}\right),$
where we have also used the fact that we are dealing with the case $\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\leq\varepsilon$. Observe that
$\overline{C}^{2}=\overline{C^{2}}-\frac{1}{\left|\Omega\right|}\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$ $=\overline{C^{2}}+\overline{A^{2}}-\overline{A^{2}}-\frac{1}{\left|\Omega\right|}\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$
$\geq M_{1}-\varepsilon\left(1+\frac{2}{\left|\Omega\right|}\right),$
thanks to the mass conservation property (4), the bound on $\overline{A^{2}}$ from above and the fact that $\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\leq\varepsilon$. Now, using the algebraic identity $(p-q)^{2}\geq\frac{p^{2}}{2}-q^{2}$, we arrive at
$\left\|\overline{A}B-\overline{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}=\frac{\left|\Omega\right|}{2}\left(\overline{C}^{2}-2\overline{A}^{2}\overline{B^{2}}\right)\geq\frac{\left|\Omega\right|}{2}\left(M_{1}-\varepsilon\left(1+\frac{2}{\left|\Omega\right|}\right)-2\varepsilon M_{2}\right)$
where we have used the aforementioned lower bound for $\overline{C}^{2}$, the mass conservation property (5) and that $\overline{A}<\sqrt{\varepsilon}$. Let us now choose
(41) $\varepsilon:=\frac{M_{1}}{2}\frac{\left|\Omega\right|}{\left|\Omega\right|+2+2M_{2}\left|\Omega\right|}.$
With the above choice of $\varepsilon$, we obtain
(42) $\left\|\overline{A}B-\overline{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\geq\frac{M_{1}\left|\Omega\right|}{4}\geq\frac{M_{1}}{4M_{2}}\left\|B-\overline{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)},$
where we have used the observation (43) from earlier. Using (40) and (42) in (39) and (36) helps us deduce that
$D(a,b,c)$ $\geq\left(\frac{4d_{a}}{P(\Omega)}-2\,\eta\,K_{3}\,(1+t)^{\frac{N-2}{N-1}}\right)\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}+\left(\frac{4d_{c}}{P(\Omega)}-2\,\eta\left|\Omega\right|^{\frac{2}{N}}\right)\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}+\eta\varepsilon\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$
for the case $\overline{A}\geq\sqrt{\varepsilon}$ and that
$D(a,b,c)$ $\geq\left(\frac{4d_{a}}{P(\Omega)}-2\,\eta\,K_{3}\,(1+t)^{\frac{N-2}{N-1}}\right)\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}+\left(\frac{4d_{c}}{P(\Omega)}-2\,\eta\left|\Omega\right|^{\frac{2}{N}}\right)\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}+\eta\frac{M_{1}}{8M_{2}}\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$
for the case $\overline{A}<\sqrt{\varepsilon}$. Next, observe that by taking
$\eta(t):=\left(\frac{2\min\{d_{a},d_{c},2\}}{P(\Omega)\left(K_{3}+\left|\Omega\right|^{\frac{2}{N}}\right)+1}\right)(1+t)^{-\frac{N-2}{N-1}}$
we obtain
$D(a,b,c)\geq\min\left\{\frac{M_{1}\left|\Omega\right|}{2\left|\Omega\right|+4+4M_{2}\left|\Omega\right|},\frac{M_{1}}{8M_{2}}\right\}\eta(t)\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}.$
Observe that the above choice of $\eta$ does satisfy $0\leq\eta\leq 4$. Hence we deduce in the case corresponding to (37) that
$D(a,b,c)$ $=\frac{1}{2}D(a,b,c)+\frac{1}{2}D(a,b,c)$
$\geq\frac{2d_{a}\left|\Omega\right|^{-\frac{2}{N}}}{P(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{2d_{c}\left|\Omega\right|^{-\frac{2}{N}}}{P(\Omega)}\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{1}{2}\min\left\{\frac{M_{1}\left|\Omega\right|}{2\left|\Omega\right|+4+4M_{2}\left|\Omega\right|},\frac{M_{1}}{8M_{2}}\right\}\eta(t)\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$
$\geq\mathcal{K}_{1}(1+t)^{-\frac{N-2}{N-1}}\left(\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\right)$
for all $t\geq 0$, where
$\mathcal{K}_{1}:=\min\left\{\frac{2d_{a}\left|\Omega\right|^{-\frac{2}{N}}}{P(\Omega)},\frac{2d_{c}\left|\Omega\right|^{-\frac{2}{N}}}{P(\Omega)},\min\left\{\frac{M_{1}\left|\Omega\right|}{2\left|\Omega\right|+4-4M_{2}\left|\Omega\right|},\frac{M_{1}}{8M_{2}}\right\}\left(\frac{\min\{d_{a},d_{c},2\}}{P(\Omega)\left(K_{3}+\left|\Omega\right|^{\frac{2}{N}}\right)+1}\right)\right\}.$
The case corresponding to (38) is relatively simpler. Observe that
$D(a,b,c)$ $\geq\frac{4\left|\Omega\right|^{-\frac{2}{N}}}{P(\Omega)}\min\left\{d_{a},d_{c}\right\}\max\left\{\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)},\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\right\}$
$>\frac{4\varepsilon\left|\Omega\right|^{-\frac{2}{N}}}{P(\Omega)}\min\left\{d_{a},d_{c}\right\}\geq\frac{4\varepsilon\left|\Omega\right|^{-\frac{2}{N}}}{P(\Omega)M_{2}\left|\Omega\right|}\min\left\{d_{a},d_{c}\right\}\left\|B-\overline{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)},$
where the final inequality is thanks to the following observation:
(43) $\left\|B-\overline{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}=\left\|B\right\|^{2}_{\mathrm{L}^{2}(\Omega)}-\overline{B}\left|\Omega\right|\leq\left\|B\right\|^{2}_{\mathrm{L}^{2}(\Omega)}=\left\|b\right\|_{\mathrm{L}^{1}(\Omega)}\leq M_{2}\left|\Omega\right|,$
which is a consequence of the mass conservation property (5). Hence we deduce in this case that
$D(a,b,c)$ $=\frac{1}{2}D(a,b,c)+\frac{1}{2}D(a,b,c)$
$\geq\frac{2d_{a}\left|\Omega\right|^{-\frac{2}{N}}}{P(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{2d_{c}\left|\Omega\right|^{-\frac{2}{N}}}{P(\Omega)}\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{2\varepsilon\left|\Omega\right|^{-\frac{2}{N}}\min\left\{d_{a},d_{c}\right\}}{P(\Omega)M_{2}\left|\Omega\right|}\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$
$\geq\mathcal{K}_{2}(1+t)^{-\frac{N-2}{N-1}}\left(\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\right)$
for all $t\geq 0$, where
$\mathcal{K}_{2}:=\min\left\{\frac{2d_{a}\left|\Omega\right|^{-\frac{2}{N}}}{P(\Omega)},\frac{2d_{c}\left|\Omega\right|^{-\frac{2}{N}}}{P(\Omega)},\frac{2\varepsilon\left|\Omega\right|^{-\frac{2}{N}}\min\left\{d_{a},d_{c}\right\}}{P(\Omega)M_{2}\left|\Omega\right|}\right\}.$
Taking $\mathcal{K}:=\min\{\mathcal{K}_{1},\mathcal{K}_{2}\}$ yields the desired result. ∎
It should be noted that a result similar to that of Proposition 4 can be found when the dimension $N<4$. As in the proof of the above proposition, employing the Poincaré-Wirtinger inequality in the expression (34) for the entropy dissipation for the case of $N<4$ yields
$D(a,b,c)\geq\frac{4d_{a}}{P(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}+\frac{4d_{c}}{P(\Omega)}\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}+4\left\|AB-C\right\|^{2}_{\mathrm{L}^{2}(\Omega)}.$
Arguing exactly as in the proof of Proposition 4 and exploiting the bound
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}\leq\hat{K}(1+t)^{\frac{5}{6}}\qquad\mbox{ for }\,t\geq 0,$
obtained in Proposition 3 and Lemma 3 helps us prove the following result. To avoid the repeat of arguments, we skip its proof. The key point to be noted, however, is that this result is unconditional in the sense that the nonzero diffusion coefficients are not assumed to satisfy the closeness condition (10).
Proposition 5.
Let $N<4$ and let $(a,b,c)$ be the solution to the degenrate system (2). The entropy dissipation $D(a,b,c)$ satisfies
$D(a,b,c)\geq\mathcal{S}\,(1+t)^{-\frac{5}{6}}\left(\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\right)\qquad\mbox{ for }t\geq 0,$
where the positive constant $\mathcal{S}$ depends only on the dimension N, the domain $\Omega$, the constants $M_{1}$ and $M_{2}$ in the mass conservation properties (4)-(5) and the nonzero diffusion coefficients $d_{a}$ and $d_{c}$.
Next, we derive a sub-exponential decay estimate for the relative entropy.
Proposition 6.
Let $N\geq 4$ and let $(a,b,c)$ be the solution to the degenerate system (2). Let $(a_{\infty},b_{\infty},c_{\infty})$ be the associated equilibrium state given by (6)-(8). Let the nonzero diffusion coefficients $d_{a},d_{c}$ satisfy the closeness condition (10). Then, for any given positive $\varepsilon\ll 1$, there exists a time $T_{\varepsilon}$ and two positive constants $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ such that
(44) $E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})\leq\mathcal{S}_{1}\,e^{-\mathcal{S}_{2}(1+t)^{\frac{1-\epsilon}{N-1}}}\qquad\mbox{ for }\,t\geq T_{\varepsilon}.$
Proof.
The relative entropy reads
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})$ $=\int_{\Omega}\left(a\ln{a}-a-a_{\infty}\ln{a_{\infty}}+a_{\infty}\right)\,{\rm d}x+\int_{\Omega}\left(b\ln{b}-b-b_{\infty}\ln{b_{\infty}}+b_{\infty}\right)\,{\rm d}x$
$\quad+\int_{\Omega}\left(c\ln{c}-c-c_{\infty}\ln{c_{\infty}}+c_{\infty}\right)\,{\rm d}x.$
Using the relation (7), the above expression for the relative entropy becomes
(45) $E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})$ $=\int_{\Omega}\left(a\ln{\frac{a}{a_{\infty}}}-a+a_{\infty}\right)\,{\rm d}x+\int_{\Omega}\left(b\ln{\frac{b}{b_{\infty}}}-b+b_{\infty}\right)\,{\rm d}x$
$\quad+\int_{\Omega}\left(c\ln{\frac{c}{c_{\infty}}}-c+c_{\infty}\right)\,{\rm d}x.$
Let us define a function $\Phi:(0,\infty)\times(0,\infty)\to\mathbb{R}$ as follows:
(46) $\Phi(x,y):=\left\{\begin{aligned} \frac{x\ln\left(\frac{x}{y}\right)-x+y}{\left(\sqrt{x}-\sqrt{y}\right)^{2}}&\qquad\mbox{ for }x\not=y,\\ 2&\qquad\mbox{ for }x=y.\end{aligned}\right.$
It can be shown (see [DF06, Lemma 2.1, p.162] for details) that the above defined function satisfies the following bound:
(47) $\Gamma(x,y)\leq C_{\Phi}\max\left\{1,\ln\left(\frac{x}{y}\right)\right\}$
for some positive constant $C_{\Phi}$. Next we rewrite the relative entropy as
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})$ $=\int_{\Omega}\Phi(a,a_{\infty})(A-A_{\infty})^{2}\,{\rm d}x+\int_{\Omega}\Phi(b,b_{\infty})(B-B_{\infty})^{2}\,{\rm d}x+\int_{\Omega}\Phi(c,c_{\infty})(C-C_{\infty})^{2}\,{\rm d}x.$
Note that for any $p\geq\frac{1}{2}$ and $q>0$, we have
$\ln p-\ln q\leq\ln(1+\left|p\right|)+\left|\ln q\right|\leq 1+\ln\left|p\right|+\left|\ln q\right|$
and for any $0<p<\frac{1}{2}$ and $q>0$, we have
$\ln p-\ln q\leq\left|\ln q\right|.$
This helps us arrive at
(48) $E(a,b,c)-$ $E(a_{\infty},b_{\infty},c_{\infty})$
$\leq C_{1}\left(1+\ln(1+t)\right)\left(\left\|A-A_{\infty}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|B-B_{\infty}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|C-C_{\infty}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\right)$
for all $t\geq 0$, where the positive constant $C_{1}$ is given by
$C_{1}:=C_{\Phi}\left(1+\left|\ln a_{\infty}\right|+\left|\ln b_{\infty}\right|+\left|\ln c_{\infty}\right|+\left|\ln K_{\infty}\right|+\mu\right).$
Here the constants $K_{\infty}$ and $\mu$ are the ones appearing in the $\mathrm{L}^{\infty}(\Omega_{t})$ bounds on the concentrations from Lemma 4. The factor $\left(1+\ln(1+t)\right)$ in the above estimate of the relative entropy is due to the fact that Lemma 4 says that at least one of the concentrations has a polynomial (in time) bound on its $\mathrm{L}^{\infty}$-norm. It should be noted that having an uniform (in time) bound on all of the concentrations gets rid of this time factor. In [DF06], the authors prove the following bound
(49) $\left\|A-A_{\infty}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|B-B_{\infty}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$ $+\left\|C-C_{\infty}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$
$\leq C_{2}\left(\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|AB-C\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\right)$
exploiting only the conservation properties (4)-(5). Furthermore, the constant $C_{2}$ in the above bound depends only on the equilibrium states $A_{\infty},B_{\infty},C_{\infty}$ and the constants $M_{1},M_{2}$ from the conservation properties (4)-(5) (see [DF06, Lemma 3.2, p.168] for precise expression for the constant). We thus arrive at the following bound for the relative entropy using (48) and (49):
(50) $E(a,b,c)-$ $E(a_{\infty},b_{\infty},c_{\infty})$
$\leq C_{3}\left(1+\ln(1+t)\right)\left(\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|AB-C\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\right)$
where the positive constant $C_{3}:=C_{1}C_{2}$. Hence, thanks to the lower bound (33) obtained in Proposition 4 and the lower bound (35), it follows from (50) that
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})\leq C_{3}\left(1+\ln(1+t)\right)\max\left\{\frac{1}{\mathcal{K}},\frac{1}{4}\right\}(1+t)^{\frac{N-2}{N-1}}D(a,b,c).$
Note that for any given positive $\varepsilon\ll 1$, there exists a time $T_{\varepsilon}$ such that
$\ln(1+t)<(1+t)^{\frac{\varepsilon}{N-1}}\qquad\mbox{ for all }t\geq T_{\varepsilon}.$
Hence we have
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})\leq C_{4}(1+t)^{\frac{N-2+\varepsilon}{N-1}}D(a,b,c)\qquad\mbox{ for all }t\geq T_{\varepsilon},$
where the constant $C_{4}:=C_{3}\max\left\{\frac{1}{\mathcal{K}},\frac{1}{4}\right\}$. Recall that we have
$\frac{{\rm d}}{{\rm d}t}\left(E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})\right)=-D(a,b,c)\qquad\mbox{ for all }t>0.$
Thus we have
$\frac{{\rm d}}{{\rm d}t}\left(E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})\right)\leq-\frac{1}{C_{4}}(1+t)^{-\frac{N-2+\varepsilon}{N-1}}\left(E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})\right)\qquad\mbox{ for all }t\geq T_{\varepsilon}.$
Integrating the above differential inequality, we obtain
$E$ $(a(t,\cdot),b(t,\cdot),c(t,\cdot))-E(a_{\infty},b_{\infty},c_{\infty})$
$\leq\left(E(a(T_{\varepsilon},\cdot),b(T_{\varepsilon},\cdot),c(T_{\varepsilon},\cdot))-E(a_{\infty},b_{\infty},c_{\infty})\right)e^{-\frac{N-1}{C_{4}(1-\varepsilon)}(1+T_{\varepsilon})^{\frac{1-\varepsilon}{N-1}}}e^{-\frac{N-1}{C_{4}(1-\varepsilon)}(1+t)^{\frac{1-\varepsilon}{N-1}}}$
$\leq\left(E(a_{0},b_{0},c_{0})-E(a_{\infty},b_{\infty},c_{\infty})\right)e^{-\frac{N-1}{C_{4}(1-\varepsilon)}}e^{-\frac{N-1}{C_{4}(1-\varepsilon)}(1+t)^{\frac{1-\varepsilon}{N-1}}}.$
We have thus proved the sub-exponential decay (44) of relative entropy with the following explicit constants:
$\mathcal{S}_{1}$ $=\left(E(a_{0},b_{0},c_{0})-E(a_{\infty},b_{\infty},c_{\infty})\right)e^{-\frac{N-1}{C_{4}(1-\varepsilon)}}$
$\mathcal{S}_{2}$ $=\frac{N-1}{C_{4}(1-\varepsilon)}$
A result similar to that of Proposition 6 can be found when the dimension $N<4$. The proof goes along similar lines and we skip it in the interest of space. The proof banks on the lower bound for the dissipation functional obtained in Proposition 5. The key point to be noted, however, is that this result is unconditional in the sense that the nonzero diffusion coefficients are not assumed to satisfy the closeness condition (10).
Proposition 7.
Let $N<4$ and let $(a,b,c)$ be the solution to the degenerate system (2). Let $(a_{\infty},b_{\infty},c_{\infty})$ be the associated equilibrium state given by (6)-(8). Then, for any given positive $\varepsilon\ll 1$, there exists a time $T_{\varepsilon}$ and two positive constants $\mathcal{S}_{3}$ and $\mathcal{S}_{4}$ such that
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})\leq\mathcal{S}_{3}\,e^{-\mathcal{S}_{4}(1+t)^{\frac{1-\epsilon}{6}}}\qquad\mbox{ for }\,t\geq T_{\varepsilon}.$
We are now equipped to prove our main result of this section.
Proof of Theorem 1.
We have already obtained sub-exponential decay (in time) of the relative entropy in Proposition 6 (for dimension $N\geq 4$) and in Proposition 7 (for dimension $N<4$). Hence the sub-exponential decay in the $\mathrm{L}^{1}$-norm is a direct consequence of the following Cziszár-Kullback-Pinsker type inequality that relates relative entropy and the $\mathrm{L}^{1}$-norm:
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})\geq\frac{\left(3+2\sqrt{2}\right)|\Omega|}{2M_{1}\left(9+2\sqrt{2}\right)}$ $\left\|a-a_{\infty}\right\|^{2}_{\mathrm{L}^{1}(\Omega)}+\frac{\left(3+2\sqrt{2}\right)|\Omega|}{2M_{1}\left(9+2\sqrt{2}\right)}\left\|b-b_{\infty}\right\|^{2}_{\mathrm{L}^{1}(\Omega)}$
$+\frac{\left(3+2\sqrt{2}\right)|\Omega|}{\left(M_{1}+M_{2}\right)\left(9+2\sqrt{2}\right)}\left\|c-c_{\infty}\right\|^{2}_{\mathrm{L}^{1}(\Omega)}.$
The proof of the above functional inequality is available in [DF06, Lemma 3.3, p.173] which exploits the conservation properties (4) and (5). ∎
3. The case of $d_{c}=0$
In this section, we shall devote our attention to the study of the degenerate model (3) which corresponds to the vanishing of the diffusion coefficient $d_{c}$. The existence of a smooth positive solution to (3) was proved in [DF15, Theorem 3.2] for all smooth initial non-negative data if the dimension $N\leq 3$. Our main objective of this section is to understand the large time behaviour of solutions to (3). Here, we choose to present our computations when the dimension $N=3$. Analogous results hold true in dimensions one and two as well. Note that we have the entropy equality
$\frac{{\rm d}}{{\rm d}t}\left(E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})\right)=-D(a,b,c)\qquad\mbox{ for all }t>0,$
where $(a,b,c)$ is the solution to the degenerate model (3) and $(a_{\infty},b_{\infty},c_{\infty})$ is the corresponding equilibrium state given by (6) and (8). The dissipation functional $D$ in the above equality is given by
$D(a,b,c)=4d_{a}\int_{\Omega}\left|\nabla A\right|^{2}\,{\rm d}x+4d_{b}\int_{\Omega}\left|\nabla B\right|^{2}\,{\rm d}x+\int_{\Omega}(ab-c)\ln\left(\frac{ab}{c}\right)\,{\rm d}x.$
We arrive at the following straightforward lower bound for the dissipation:
(51) $D(a,b,c)\geq\frac{4d_{a}}{P(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}+\frac{4d_{b}}{P(\Omega)}\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}+4\left\|AB-C\right\|^{2}_{\mathrm{L}^{2}(\Omega)},$
thanks to the Poincaré-Wirtinger inequality and an algebraic identity which says that for all $p,q\geq 0$, there holds $(p-q)\left(\ln p-\ln q\right)\geq 4\left(\sqrt{p}-\sqrt{q}\right)^{2}$. An application of the Hölder inequality leads to
(52) $D(a,b,c)\geq\frac{4d_{a}\left|\Omega\right|^{-\frac{2}{3}}}{P(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{4d_{b}\left|\Omega\right|^{-\frac{2}{3}}}{P(\Omega)}\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+4\left\|AB-C\right\|^{2}_{\mathrm{L}^{2}(\Omega)}.$
It is apparent from the above inequality that the term involving $\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$ is missing from its lower bound. A similar scenario was handled in the previous section while dealing with the missing $\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$ term. Our strategy was to derive polynomial (in time) bounds on the supremum norms of the concentrations. Here too, we will adapt a similar approach. Note, however, that the proofs of most results in this section markedly differ from the proofs in the previous section. We begin with a time-dependent $\mathrm{L}^{1}(\Omega)$ estimate on certain combinations of the concentrations. This result is inspired by [DFPV07, Theorem 3.1, p.495]. Our proof argues along similar lines as in [DFPV07] while keeping track of the polynomial (in time) bound.
Lemma 5.
Let $N\leq 3$ and let $(a,b,c)$ be the solution to the degenerate system (3). Then, there exist positive constants $\ell_{1}$ and $\ell_{2}$ such that for all $t\geq 0$, we have
$\int_{0}^{t}\int_{\Omega}\Big{(}a^{2}(s,x)+a(s,x)c(s,x)\Big{)}\,{\rm d}x\,{\rm d}s$ $\leq\ell_{1}\left(1+t\right),$
$\int_{0}^{t}\int_{\Omega}\Big{(}b^{2}(s,x)+b(s,x)c(s,x)\Big{)}\,{\rm d}x\,{\rm d}s$ $\leq\ell_{2}\left(1+t\right).$
Proof.
Adding up the equations for $a$ and $c$ in the degenerate model (3), we obtain
(53) $\partial_{t}\left(a+c\right)-d_{a}\Delta a=0.$
Let $H\in\mathrm{C}^{\infty}_{c}(\Omega)$ be arbitrary and let $Z:=\frac{a\,d_{a}}{a+c}$. Now consider the following backward parabolic problem:
(54) $-\partial_{t}w-Z\Delta w&=H\,\sqrt{Z}\qquad\mbox{ in }\Omega_{T},\\ \nabla w\cdot n(x)&=0\qquad\mbox{ on }\partial\Omega_{T},\\ w(T,x)&=0\qquad\mbox{ for }x\in\Omega.\end{aligned}\right.$
In this duality approach, the idea is to multiply the equation (53) by the solution $w$ to the backward problem (54) followed by an integration over $\Omega_{T}$ leading to
$-\int_{\Omega}w(0,x)\left(a_{0}(x)+c_{0}(x)\right)\,{\rm d}x-\int_{0}^{T}\int_{\Omega}\left(a+c\right)\partial_{t}w\,{\rm d}x\,{\rm d}t-\int_{0}^{T}\int_{\Omega}d_{a}a\Delta w\,{\rm d}x\,{\rm d}t=0,$
thanks to integration by parts. Using the equation satisfied by $w$ in (54), we deduce the following:
(55) $\int_{0}^{T}\int_{\Omega}\left(a+c\right)\,H\,\sqrt{Z}\,{\rm d}x\,{\rm d}t=\int_{\Omega}w(0,x)\left(a_{0}(x)+c_{0}(x)\right)\,{\rm d}x\leq\left\|w(0,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}\left\|a_{0}+c_{0}\right\|_{\mathrm{L}^{2}(\Omega)},$
where the inequality is due to the Cauchy-Schwarz inequality. Our next objective is to get an estimate for $\left\|w(0,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}$. To that end, multiply the evolution equation by $-\Delta w$ and integrating with respect to the spatial variable yields
$\int_{\Omega}\Delta w\partial_{t}w\,{\rm d}x+\int_{\Omega}Z\left(\Delta w\right)^{2}\,{\rm d}x=-\int_{\Omega}H\sqrt{Z}\Delta w\,{\rm d}x.$
Performing an integration by parts in the first term on the left hand side (while using the homogeneous Neumann boundary condition from (54)) and employing Young’s inequality for the term on the right hand side yields
$-\frac{1}{2}\frac{{\rm d}}{{\rm d}t}\int_{\Omega}\left|\nabla w\right|^{2}\,{\rm d}x+\int_{\Omega}Z\left(\Delta w\right)^{2}\,{\rm d}x\leq\frac{1}{2}\int_{\Omega}H^{2}\,{\rm d}x+\frac{1}{2}\int_{\Omega}Z\left(\Delta w\right)^{2}\,{\rm d}x.$
Integrating the above inequality on the interval $(0,T)$ in the time variable results in
(56) $\int_{\Omega}\left|\nabla w(0,x)\right|^{2}\,{\rm d}x+\int_{0}^{T}\int_{\Omega}Z\left(\Delta w\right)^{2}\,{\rm d}x\,{\rm d}t\leq\int_{0}^{T}\int_{\Omega}H^{2}\,{\rm d}x\,{\rm d}t.$
Hence, invoking Poincaré-Wirtinger inequality, we get
$\int_{\Omega}\left(w(0,x)-\frac{1}{\left|\Omega\right|}\int_{\Omega}w(0,y)\,{\rm d}y\right)^{2}\,{\rm d}x\leq P(\Omega)\int_{\Omega}\left|\nabla w(0,x)\right|^{2}\,{\rm d}x,$
where $P(\Omega)$ is the Poincaré constant. Note that integrating the evolution equation in (54) over $\Omega_{T}$ yields
$\int_{\Omega}w(0,x)\,{\rm d}x$ $=\int_{0}^{T}\int_{\Omega}Z\Delta w\,{\rm d}x\,{\rm d}t+\int_{0}^{T}\int_{\Omega}H\sqrt{Z}\,{\rm d}x\,{\rm d}t$
$\leq\left(\left(\int_{0}^{T}\int_{\Omega}Z\left(\Delta w\right)^{2}\,{\rm d}x\,{\rm d}t\right)^{\frac{1}{2}}+\left(\int_{0}^{T}\int_{\Omega}H^{2}\,{\rm d}x\,{\rm d}t\right)^{\frac{1}{2}}\right)\left\|\sqrt{Z}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\leq 2\left\|H\right\|_{\mathrm{L}^{2}(\Omega_{T})}\left\|\sqrt{Z}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
thanks to the estimate from (56). Using the fact that $Z\leq d_{a}$, we deduce
(57) $\left(\int_{\Omega}w(0,x)\,{\rm d}x\right)^{2}\leq 4d_{a}T\left|\Omega\right|\left\|H\right\|^{2}_{\mathrm{L}^{2}(\Omega_{T})}.$
Hence we obtain
$\int_{\Omega}\left|w(0,x)\right|^{2}\,{\rm d}x\leq 2\int_{\Omega}\left(w(0,x)-\frac{1}{\left|\Omega\right|}\int_{\Omega}w(0,y)\,{\rm d}y\right)^{2}\,{\rm d}x+\frac{2}{\left|\Omega\right|}\left(\int_{\Omega}w(0,x)\,{\rm d}x\right)^{2}\leq\left(2P(\Omega)+8d_{a}T\right)\left\|H\right\|^{2}_{\mathrm{L}^{2}(\Omega_{T})},$
thanks to (56) and (57). Going back to (55), we have thus obtained
$\int_{0}^{T}\int_{\Omega}\left(a+c\right)\,H\,\sqrt{Z}\,{\rm d}x\,{\rm d}t\leq\left(2P(\Omega)+8d_{a}T\right)^{\frac{1}{2}}\left\|a_{0}+c_{0}\right\|_{\mathrm{L}^{2}(\Omega)}\left\|H\right\|_{\mathrm{L}^{2}(\Omega_{T})}.$
Since the above inequality holds true for arbitrary $H\in\mathrm{C}^{\infty}_{c}(\Omega)$, we deduce by duality that
$\int_{0}^{T}\int_{\Omega}\left(a+c\right)^{2}Z\,{\rm d}x\,{\rm d}t\leq\left(2P(\Omega)+8d_{a}T\right)\left\|a_{0}+c_{0}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$
Substituting for $Z$ in the above inequality, we arrive at
$\int_{0}^{T}\int_{\Omega}\left(a^{2}+ac\right)\,{\rm d}x\,{\rm d}t\leq\left(\frac{2P(\Omega)+8d_{a}}{d_{a}}\right)\left\|a_{0}+c_{0}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}(1+T).$
Proceeding exactly as above but working with the equation satisfied by $b+c$, we can obtain
$\int_{0}^{T}\int_{\Omega}\left(b^{2}+bc\right)\,{\rm d}x\,{\rm d}t\leq\left(\frac{2P(\Omega)+8d_{b}}{d_{b}}\right)\left\|b_{0}+c_{0}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}(1+T).$
This concludes the proof. ∎
Lemma 6.
Let $N\leq 3$ and let $(a,b,c)$ be the solution to the degenerate system (3). Then, there exist positive constants $\ell_{3},\ell_{4},\ell_{7}$ such that
$\left\|a(t,\cdot)\right\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}$ $\leq\ell_{3}\left(1+t\right)^{\frac{1}{3}}\qquad\mbox{ for }\,t\geq 0,$
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}$ $\leq\ell_{4}\left(1+t\right)^{\frac{1}{3}}\qquad\mbox{ for }\,t\geq 0,$
$\left\|c(t,\cdot)\right\|_{\mathrm{L}^{3}(\Omega)}$ $\leq\ell_{7}\left(1+t\right)\qquad\,\,\mbox{ for }\,t\geq 0.$
Proof.
Multiplying the equation for $a$ in (3) by $a$ and integrating in space and time variables yields
$\frac{1}{2}\int_{\Omega}a^{2}\,{\rm d}x+d_{a}\int_{0}^{t}\int_{\Omega}\left|\nabla a\right|^{2}\,{\rm d}x\,{\rm d}s\leq\frac{1}{2}\int_{\Omega}a^{2}_{0}\,{\rm d}x+\int_{0}^{t}\int_{\Omega}ac\,{\rm d}x\,{\rm d}s,$
where we have used the fact that $a,b$ are non-negative. The estimate from Lemma 5 helps us get
(58) $\left\|a(t,\cdot)\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+d_{a}\left\|\nabla a\right\|^{2}_{\mathrm{L}^{2}(\Omega_{t})}\leq\left\|a_{0}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+2\ell_{1}\left(1+t\right).$
As
$\frac{\frac{1}{3}}{1}+\frac{1-\frac{1}{3}}{2}=\frac{2}{3},$
interpolation leads to the following bound:
$\left\|a(t,\cdot)\right\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}\leq\left\|a(t,\cdot)\right\|^{\frac{1}{3}}_{\mathrm{L}^{1}(\Omega)}\left\|a(t,\cdot)\right\|^{\frac{2}{3}}_{\mathrm{L}^{2}(\Omega)}\leq M_{1}^{\frac{1}{3}}\left|\Omega\right|^{\frac{1}{3}}\left(\left\|a_{0}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+2\ell_{1}\left(1+t\right)\right)^{\frac{1}{3}}$
where we have used the mass conservation property (4). Taking $\ell_{3}:=\left(M_{1}\left|\Omega\right|\left(\left\|a_{0}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+2\ell_{1}\right)\right)^{\frac{1}{3}}$, we have thus shown
$\left\|a(t,\cdot)\right\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}\leq\ell_{3}\left(1+t\right)^{\frac{1}{3}}\qquad\mbox{ for }\,t\geq 0.$
Arguing exactly as above, we obtain
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}\leq\ell_{4}\left(1+t\right)^{\frac{1}{3}}\qquad\mbox{ for }\,t\geq 0,$
with the constant $\ell_{4}=\left(M_{2}\left|\Omega\right|\left(\left\|b_{0}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+2\ell_{2}\right)\right)^{\frac{1}{3}}$. It follows from (58) and the estimate from Lemma 5 that
(59) $\int_{0}^{t}\left\|a(s,\cdot)\right\|^{2}_{\mathrm{H}^{1}(\Omega)}\,{\rm d}s\leq\ell_{5}(1+t).$
A similar estimate holds for $\left\|b\right\|^{2}_{\mathrm{L}^{2}(0,t;\mathrm{H}^{1}(\Omega))}$ as well. By Sobolev embedding we have
(60) $\left\|a(t,\cdot)\right\|_{\mathrm{L}^{6}(\Omega)}\leq\ell_{6}\left\|a(t,\cdot)\right\|_{\mathrm{H}^{1}(\Omega)},\qquad\left\|b(t,\cdot)\right\|_{\mathrm{L}^{6}(\Omega)}\leq\ell_{6}\left\|b(t,\cdot)\right\|_{\mathrm{H}^{1}(\Omega)}$
Exploiting the non-negativity of $c$, observe from (3) that $c$ satisfies the inequality $\partial_{t}c\leq ab$. Hence we have
$\left(c(t,x)\right)^{3}\leq 2^{2}\left(\left(c_{0}(x)\right)^{3}+\left(\int_{0}^{t}a(s,x)\,b(s,x)\,{\rm d}s\right)^{3}\right)\leq 2^{2}(1+t)^{2}\left(\left(c_{0}(x)\right)^{3}+\int_{0}^{t}\left(a(s,x)\,b(s,x)\right)^{3}\,{\rm d}s\right),$
thanks to Jensen’s inequality. Integrating the above inequality in the $x$ variable yields
$\left\|c(t,\cdot)\right\|^{3}_{\mathrm{L}^{3}(\Omega)}\leq 2^{2}(1+t)^{2}\left(\left\|c_{0}\right\|^{3}_{\mathrm{L}^{3}(\Omega)}+\int_{0}^{t}\left\|ab(s,\cdot)\right\|^{3}_{\mathrm{L}^{3}(\Omega)}\,{\rm d}s\right)$
Employing the Hölder inequality leads to the following bound
$\left\|c(t,\cdot)\right\|^{3}_{\mathrm{L}^{3}(\Omega)}$ $\leq 2^{2}(1+t)^{2}\left(\left\|c_{0}\right\|^{3}_{\mathrm{L}^{3}(\Omega)}+\int_{0}^{t}\left\|a(s,\cdot)\right\|^{\frac{1}{2}}_{\mathrm{L}^{6}(\Omega)}\,\left\|b(s,\cdot)\right\|^{\frac{1}{2}}_{\mathrm{L}^{6}(\Omega)}\,{\rm d}s\right)$
$\leq 2^{2}(1+t)^{2}\left(\left\|c_{0}\right\|^{3}_{\mathrm{L}^{3}(\Omega)}+\frac{1}{2}\int_{0}^{t}\left(\left\|a(s,\cdot)\right\|_{\mathrm{L}^{6}(\Omega)}+\left\|b(s,\cdot)\right\|_{\mathrm{L}^{6}(\Omega)}\right)\,{\rm d}s\right),$
where we have applied the Young’s inequality. Hence it follows from (59) and (60) that
$\left\|c(t,\cdot)\right\|_{\mathrm{L}^{3}(\Omega)}\leq\ell_{7}(1+t)\qquad\mbox{ for }t\geq 0,$
for some constant $\ell_{7}$
Our next task is to obtain polynomial (in time) growth estimates on the solution in the supremum norm.
Proposition 8.
Let $N\leq 3$ and let $(a,b,c)$ be the solution to the degenerate system (3). Then, there exist positive constants $K_{\infty}$ and $\mu$ such that for all $t\geq 0$, we have
$\left\|a\right\|_{\mathrm{L}^{\infty}(\Omega_{t})}$ $\leq K_{\infty}\left(1+t\right)^{\mu}$
$\left\|b\right\|_{\mathrm{L}^{\infty}(\Omega_{t})}$ $\leq K_{\infty}\left(1+t\right)^{\mu}$
$\left\|c\right\|_{\mathrm{L}^{\infty}(\Omega_{t})}$ $\leq K_{\infty}\left(1+t\right)^{\mu}$
Proof.
Let $G_{d_{a}}$ denote the Green’s function associated with the operator $\partial_{t}-d_{a}\Delta$ with Neumann boundary condition. We can express the solution $a$ as follows:
(61) $a(t,x)=\tilde{a}(t,x)+\int_{0}^{t}\int_{\Omega}G_{d_{a}}(t-s,x,y)\left[c-ab\right](s,y)\,{\rm d}y\,{\rm d}s,$
where $\tilde{a}$ solves the following initial boundary value problem:
$\partial_{t}\tilde{a}-d_{a}\Delta\tilde{a}&=0\qquad\mbox{ for }t>0,\,x\in\Omega,\\ \nabla\tilde{a}\cdot n(x)&=0\qquad\mbox{ for }t>0,\,x\in\partial\Omega,\\ \tilde{a}(0,x)&=a_{0}\qquad\mbox{ for }x\in\Omega.\end{aligned}\right.$
We recall the following Gaussian bound on the Neumann Green’s function (see [Mor83, Theorem 2.2, p.37]): there exist positive constants $\texttt{C}_{H},\kappa$ such that
(62) $\left|G_{d_{a}}(t-s,x,y)\right|\leq\texttt{C}_{H}\frac{1}{(t-s)^{\frac{N}{2}}}e^{-\kappa\frac{\left|x-y\right|^{2}}{(t-s)}}=:g(t-s,x-y)$
Also see [CK15, Theorem 3.1, p.639] for general parabolic operators. As a consequence we have the following bound on the solution $\tilde{a}$ to the above homogeneous problem:
(63) $\left\|\tilde{a}(t,\cdot)\right\|_{\mathrm{L}^{p}(\Omega)}\leq C_{S}\left\|a_{0}\right\|_{\mathrm{L}^{p}(\Omega)}$
for some positive constant $C_{S}$, independent of time, and for any $p\geq 1$. In (61), the positivity of $a$ and $b$ leads to
$a(t,x)\leq\tilde{a}(t,x)+\int_{0}^{t}\int_{\Omega}G_{d_{a}}(t-s,x,y)c(s,y)\,{\rm d}y\,{\rm d}s.$
Using the aforementioned Gaussian bound, we arrive at
$a(t,x)\leq\tilde{a}(t,x)+\int_{0}^{t}\int_{\Omega}g(t-s,x-y)c(s,y)\,{\rm d}y\,{\rm d}s.$
Computing the $\mathrm{L}^{p}$ norm in the $x$ variable, the above inequality leads to
$\left\|a(t,\cdot)\right\|_{\mathrm{L}^{p}(\Omega)}\leq C_{S}\left\|a_{0}\right\|_{\mathrm{L}^{p}(\Omega)}+\int_{0}^{t}\left\|g(t-s,\cdot)\right\|_{\mathrm{L}^{r}(\Omega)}\left\|c(s,\cdot)\right\|_{\mathrm{L}^{q}(\Omega)}\,{\rm d}s,$
thanks to the bound (63), the Minkowski’s integral inequality and the Young’s convolution inequality with
$1+\frac{1}{p}=\frac{1}{r}+\frac{1}{q}.$
Therefore, there exists a positive constant $C_{H,N,r}$ such that
(64) $\left\|a(t,\cdot)\right\|_{\mathrm{L}^{p}(\Omega)}\leq C_{S}\left\|a_{0}\right\|_{\mathrm{L}^{p}(\Omega)}+C_{H,N,r}\int_{0}^{t}(t-s)^{-\frac{N}{2}\left(\frac{1}{q}-\frac{1}{p}\right)}\left\|c(s,\cdot)\right\|_{\mathrm{L}^{q}(\Omega)}\,{\rm d}s$
Taking $q=3$ and $p=\infty$ in the above bound, we obtain
$\left\|a(t,\cdot)\right\|_{\mathrm{L}^{\infty}(\Omega)}$ $\leq C_{S}\left\|a_{0}\right\|_{\mathrm{L}^{p}(\Omega)}+C_{H,N,r}\int_{0}^{t}(t-s)^{-\frac{N}{6}}\left\|c(s,\cdot)\right\|_{\mathrm{L}^{3}(\Omega)}\,{\rm d}s$
$\leq C_{S}\left\|a_{0}\right\|_{\mathrm{L}^{p}(\Omega)}+C_{H,N,r}\int_{0}^{t}(t-s)^{-\frac{N}{6}}(1+s)\,{\rm d}s,$
thanks to the bound from Lemma 6. Hence we arrive at
$\left\|a\right\|_{\mathrm{L}^{\infty}(\Omega_{t})}\leq K_{\infty}(1+t)^{\frac{18-N}{6}}$
for some positive constant $K_{\infty}$. Arguing along exactly same lines, we can obtain an estimate of $b$ in the supremum norm as well. Again, exploiting the positivity of $c$, we have from (3)
$c(t,x)\leq c_{0}(x)+\int_{0}^{t}a(s,x)b(s,x)\,{\rm d}s.$
The above supremum norm estimates on $a$ and $b$ will help us arrive at the supremum norm estimate for $c$ as well. ∎
Recall from the lower bound in (52) that the term involving $\left\|\delta_{C}\right\|_{\mathrm{L}^{2}(\Omega)}$ is apparently missing. Similar to Propositions 4 and 5, we now derive a lower bound for the dissipation functional involving this missing term.
Proposition 9.
Let $N\leq 3$ and let $(a,b,c)$ be the solution to the degenerate system (3). Then the entropy dissipation $D(a,b,c)$ satisfies
(65) $D(a,b,c)\geq\mathcal{K}_{c}\,(1+t)^{-\frac{1}{3}}\left(\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\right)\qquad\mbox{ for }t\geq 0,$
where the positive constant $\mathcal{K}_{c}$ depends only on the domain $\Omega$, the constants $M_{1}$ and $M_{2}$ in the mass conservation properties (4)-(5) and the nonzero diffusion coefficients $d_{a}$ and $d_{b}$.
Proof.
It follows from (51) that
(66) $D(a,b,c)\geq\frac{4d_{a}}{P(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}+\frac{4d_{b}}{P(\Omega)}\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}+\eta\left\|AB-C\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$
for any $0\leq\eta\leq 4$. In order to relate the dissipation functional to the missing $\left\|\delta_{C}\right\|_{\mathrm{L}^{2}(\Omega)}$ term, we work on the following term:
$\left\|AB-C\right\|_{\mathrm{L}^{2}(\Omega)}.$
Note that we have
$\left\|AB-C\right\|_{\mathrm{L}^{2}(\Omega)}$ $=\left\|\left(\delta_{A}+\overline{A}\right)\left(\delta_{B}+\overline{B}\right)-C\right\|_{\mathrm{L}^{2}(\Omega)}$
$\geq\frac{1}{2}\left\|\overline{A}\,\overline{B}-C\right\|_{\mathrm{L}^{2}(\Omega)}-3\left\|\overline{A}\,\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}-3\left\|\overline{B}\,\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}-3\left\|\delta_{A}\,\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)},$
where we have used the following algebraic identities that hold for all $p,q,r\in\mathbb{R}$:
$\left(p-q\right)^{2}\geq\frac{p^{2}}{2}-q^{2}\qquad\mbox{ and }\qquad\left(p+q+r\right)^{2}\leq 3\left(p^{2}+q^{2}+r^{2}\right).$
Using the mass conservation property (4), we have
$\int_{\Omega}A(t,x)\,{\rm d}x\leq\left|\Omega\right|^{\frac{1}{2}}\left(\int_{\Omega}a(t,x)\,{\rm d}x\right)^{\frac{1}{2}}\leq\left|\Omega\right|^{\frac{1}{2}}M_{1}^{\frac{1}{2}}\left|\Omega\right|^{\frac{1}{2}}\implies\overline{A}\leq M_{1}^{\frac{1}{2}}.$
Similarly, we have $\overline{B}\leq M_{2}^{\frac{1}{2}}$. Hence we arrive at
$\left\|AB-C\right\|_{\mathrm{L}^{2}(\Omega)}\geq\frac{1}{2}\left\|\overline{A}\,\overline{B}-C\right\|_{\mathrm{L}^{2}(\Omega)}-3M_{1}^{\frac{1}{2}}\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}-3M_{2}^{\frac{1}{2}}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}-3\left\|\delta_{A}\,\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}.$
Employing the Hölder inequality in the last three terms in the above lower bound leads to
$\left\|AB-C\right\|_{\mathrm{L}^{2}(\Omega)}\geq\frac{1}{2}\left\|\overline{A}\,\overline{B}-C\right\|_{\mathrm{L}^{2}(\Omega)}-3M_{1}^{\frac{1}{2}}\left|\Omega\right|^{\frac{2}{3}}\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}-3M_{2}^{\frac{1}{2}}\left|\Omega\right|^{\frac{2}{3}}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}-3\left\|\delta_{A}^{2}\right\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}$
Using the algebraic identity $(p-q)^{2}\leq 2(p^{2}+q^{2})$ in the last term of the above lower bound, we obtain
$\left\|AB-C\right\|_{\mathrm{L}^{2}(\Omega)}\geq\frac{1}{2}\left\|\overline{A}\,\overline{B}-C\right\|_{\mathrm{L}^{2}(\Omega)}$ $-3M_{1}^{\frac{1}{2}}\left|\Omega\right|^{\frac{2}{3}}\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}-3M_{2}^{\frac{1}{2}}\left|\Omega\right|^{\frac{2}{3}}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}$
$-6\left\|A^{2}+\overline{A}^{2}\right\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}.$
By employing triangular inequality in the last term of the above lower bound, we arrive at
$\left\|AB-C\right\|_{\mathrm{L}^{2}(\Omega)}$ $\geq\frac{1}{2}\left\|\overline{A}\,\overline{B}-C\right\|_{\mathrm{L}^{2}(\Omega)}-\left|\Omega\right|^{\frac{2}{3}}\left(3M_{1}^{\frac{1}{2}}+6M_{1}\right)\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}-3M_{2}^{\frac{1}{2}}\left|\Omega\right|^{\frac{2}{3}}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}$
$\qquad\qquad\qquad\qquad-6\left\|a\right\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}$
$\geq\frac{1}{2}\left\|\overline{A}\,\overline{B}-C\right\|_{\mathrm{L}^{2}(\Omega)}-\left(\left|\Omega\right|^{\frac{2}{3}}\left(3M_{1}^{\frac{1}{2}}+6M_{1}\right)+6\ell_{3}\left(1+t\right)^{\frac{1}{3}}\right)\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}-3M_{2}^{\frac{1}{2}}\left|\Omega\right|^{\frac{2}{3}}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{6}(\Omega)},$
thanks to the estimate on $\left\|a\right\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}$ from Lemma 6. Now we claim that
$\left\|\overline{A}\,\overline{B}-C\right\|_{\mathrm{L}^{2}(\Omega)}\geq\left\|C-\overline{C}\right\|_{\mathrm{L}^{2}(\Omega)}.$
To see this, factorising $C$ as $\overline{A}\,\overline{B}(1+\mu(x))$, we get
$\left\|\overline{A}\,\overline{B}-C\right\|^{2}_{\mathrm{L}^{2}(\Omega)}=\overline{A}^{2}\,\overline{B}^{2}\,\overline{\mu^{2}}\,\left|\Omega\right|\qquad\mbox{ and }\qquad\left\|C-\overline{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}=\overline{A}^{2}\,\overline{B}^{2}\left\|\mu-\overline{\mu}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\leq\overline{A}^{2}\,\overline{B}^{2}\,\overline{\mu^{2}}\,\left|\Omega\right|.$
Putting it all together, we arrive at
$\left\|AB-C\right\|_{\mathrm{L}^{2}(\Omega)}\geq\frac{1}{2}\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}-\left(\left|\Omega\right|^{\frac{2}{3}}\left(3M_{1}^{\frac{1}{2}}+6M_{1}\right)+6\ell_{3}\left(1+t\right)^{\frac{1}{3}}\right)\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}-3M_{2}^{\frac{1}{2}}\left|\Omega\right|^{\frac{2}{3}}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}.$
Hence it follows from (66) that
(67) $D(a,b,c)\geq\left(\frac{4d_{a}}{P(\Omega)}-3\eta\,M_{2}^{\frac{1}{2}}\left|\Omega\right|^{\frac{2}{3}}\right)$ $\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}+\frac{\eta}{2}\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$
$+\left(\frac{4d_{b}}{P(\Omega)}-\eta\,\left(\left|\Omega\right|^{\frac{2}{3}}\left(3M_{1}^{\frac{1}{2}}+6M_{1}\right)+6\ell_{3}\left(1+t\right)^{\frac{1}{3}}\right)\right)\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{6}(\Omega)}.$
Let us take
$\eta(t):=\frac{4\min\left\{d_{a},d_{b},1\right\}}{P(\Omega)\left(3M_{2}^{\frac{1}{2}}\left|\Omega\right|^{\frac{2}{3}}+\left|\Omega\right|^{\frac{2}{3}}\left(3M_{1}^{\frac{1}{2}}+6M_{1}\right)+6\ell_{3}\right)+1}(1+t)^{-\frac{1}{3}}.$
Then, it follows from (67) that the dissipation functional has the following lower bound:
(68) $D(a,b,c)\geq\frac{1}{2}\eta(t)\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}.$
Observe that the above choice of $\eta$ clearly satisfies $0\leq\eta\leq 4$. To conclude our proof, we write
$D(a,b,c)$ $=\frac{1}{2}D(a,b,c)+\frac{1}{2}D(a,b,c)$
$\geq\frac{2d_{a}\left|\Omega\right|^{-\frac{2}{3}}}{P(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{2d_{b}\left|\Omega\right|^{-\frac{2}{3}}}{P(\Omega)}\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{1}{4}\eta(t)\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)},$
where we have used the lower bound from (66) for the first term while the lower bound from (68) for the second term. Hence we have proved (65) with the constant $\mathcal{K}_{c}$ given by
$\mathcal{K}_{c}:=\min\left\{\frac{2d_{a}\left|\Omega\right|^{-\frac{2}{3}}}{P(\Omega)},\frac{2d_{b}\left|\Omega\right|^{-\frac{2}{3}}}{P(\Omega)},\frac{\min\left\{d_{a},d_{b},1\right\}}{P(\Omega)\left(3M_{2}^{\frac{1}{2}}\left|\Omega\right|^{\frac{2}{3}}+\left|\Omega\right|^{\frac{2}{3}}\left(3M_{1}^{\frac{1}{2}}+6M_{1}\right)+6\ell_{3}\right)+1}\right\}.$
Thanks to the lower bound in (65), we can derive a sub-exponential decay estimate for the relative entropy.
Proposition 10.
Let $N\leq 3$ and let $(a,b,c)$ be the solution to the degenerate system (3). Let $(a_{\infty},b_{\infty},c_{\infty})$ be the associated equilibrium state given by (6)-(8). Then, for any given positive $\varepsilon\ll 1$, there exists a time $T_{\varepsilon}$ and two positive constants $\mathcal{S}_{5}$ and $\mathcal{S}_{6}$ such that
(69) $E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})\leq\mathcal{S}_{5}\,e^{-\mathcal{S}_{6}(1+t)^{\frac{2-\epsilon}{3}}}\qquad\mbox{ for }\,t\geq T_{\varepsilon}.$
The proof of the above proposition is exactly similar to the proof of Proposition 6. Hence we skip the proof. Note that the constants $\mathcal{S}_{5}$ and $\mathcal{S}_{6}$ appearing in (69) depend on the constant $\mathcal{K}_{c}$ (appearing in Proposition 65), the constants $K_{\infty}$ and $\mu$ (appearing in Proposition 8). Finally, we are all equipped to prove our main result of this section.
Proof of Theorem 2.
We have already obtained sub-exponential decay (in time) of the relative entropy (69) in Proposition 10. Hence the sub-exponential decay in the $\mathrm{L}^{1}$-norm is a direct consequence of the following Cziszár-Kullback-Pinsker type inequality that relates relative entropy and the $\mathrm{L}^{1}$-norm:
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})\geq\frac{\left(3+2\sqrt{2}\right)|\Omega|}{2M_{1}\left(9+2\sqrt{2}\right)}$ $\left\|a-a_{\infty}\right\|^{2}_{\mathrm{L}^{1}(\Omega)}+\frac{\left(3+2\sqrt{2}\right)|\Omega|}{2M_{1}\left(9+2\sqrt{2}\right)}\left\|b-b_{\infty}\right\|^{2}_{\mathrm{L}^{1}(\Omega)}$
$+\frac{\left(3+2\sqrt{2}\right)|\Omega|}{\left(M_{1}+M_{2}\right)\left(9+2\sqrt{2}\right)}\left\|c-c_{\infty}\right\|^{2}_{\mathrm{L}^{1}(\Omega)}.$
The proof of the above functional inequality is available in [DF06, Lemma 3.3, p.173] which exploits the conservation properties (4) and (5). ∎
Appendix A some useful results
Lemma 7 (Poincaré-Wirtinger inequality).
There exists a positive constant $P(\Omega)$, depending only on $\Omega$ and $q$, such that
$P(\Omega)\left\|\nabla f\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\geq\left\|f-\overline{f}\right\|^{2}_{\mathrm{L}^{q}(\Omega)}\qquad\mbox{ for all }f\in\mathrm{H}^{1}(\Omega),$
where
$q=\left\{\begin{aligned} \frac{2N}{N-2}&\qquad\mbox{ for }\,\,N\geq 3,\\ \in[1,\infty)&\qquad\mbox{ for }\,\,N=2,\\ \in[1,\infty)\cup\{\infty\}&\qquad\mbox{ for }\,\,N=1.\end{aligned}\right.$
We refer to $P(\Omega)$ as the Poincaré constant.
Theorem 11 (Second Order Regularity and Integrability estimation).
Let $d>0$ and let $\tau\in[0,T)$. Take $\theta\in\mathrm{L}^{p}(\Omega_{\tau,T})$ for some $1<p<+\infty$. Let $\psi$ be the solution to the backward heat equation:
$\partial_{t}\psi+d\Delta\psi&={\color[rgb]{1,0,0}-\theta}\qquad\mbox{ for }(t,x)\in\Omega_{\tau,T},\\ \nabla\psi\cdot n(x)&=0\qquad\mbox{ for }(t,x)\in[\tau,T]\times\partial\Omega,\\ \psi(T,x)&=0\qquad\mbox{ for }x\in\Omega.\end{aligned}\right.$
Then, there exists a positive constant $C_{SOR}$, depending only on the domain $\Omega$, the dimension $N$ and the exponent $p$ such that the following maximal regularity holds:
(70) $\|\Delta\psi\|_{\mathrm{L}^{p}(\Omega_{\tau,T})}\leq\frac{C_{SOR}}{d}\|\theta\|_{\mathrm{L}^{p}(\Omega_{\tau,T})}.$
Moreover, if  $\theta\geq 0$ then $\psi(t,x)\geq 0$ for almost every $(t,x)\in\Omega_{\tau,T}$. Furthermore, we have
(71) If $p<\frac{N+2}{2}\,\mbox{ then }\qquad\left\|\psi\right\|_{\mathrm{L}^{s}(\Omega_{\tau,T})}\leq C_{IE}\left\|\theta\right\|_{\mathrm{L}^{p}(\Omega_{\tau,T})}\qquad\mbox{ for all }\,\,s<\frac{(N+2)p}{N+2-2p}$
If $p=\frac{N+2}{2}\,\mbox{ then }\qquad\left\|\psi\right\|_{\mathrm{L}^{s}(\Omega_{\tau,T})}\leq C_{IE}\left\|\theta\right\|_{\mathrm{L}^{p}(\Omega_{\tau,T})}\qquad\mbox{ for all }\,\,s<\infty$
where the constant $C_{IE}=C_{IE}(T-\tau,\Omega,d,p,s)$ and
(72) if $p>\frac{N+2}{2}\,\mbox{ then }\qquad\left\|\psi\right\|_{\mathrm{L}^{\infty}(\Omega_{\tau,T})}\leq C_{IE}\left\|\theta\right\|_{\mathrm{L}^{p}(\Omega_{\tau,T})}.$
The proof of (70) can be found in [Lam87, Theorem 1]. We refer to the constant $C_{SOR}$ as the second order regularity constant. Proof of the estimates (71) can be found in [CDF14, Lemma 3.3] and the estimate (72) was derived in [Tan18, Lemma 4.6]. We refer to the constant $C_{IE}$ as the integrability estimation constant.
Theorem 12 ($p^{\textrm{th}}$ order integrability estimation).
Let $p\in(2,\infty)$ and let $p^{\prime}$ be its Hölder conjugate. Let $M(t,x)$ be such that the following holds
$\theta\leq M(t,x)\leq\Theta\quad\forall(t,x)\in\Omega_{T},$
for some fixed positive constants $\theta,\Theta$. Let $\psi_{0}\in\mathrm{L}^{p}(\Omega)$ and let $\psi$ be a weak solution to
$\partial_{t}\psi-\Delta\left(M\psi\right)&=0\qquad\qquad\mbox{ for }(t,x)\in\Omega_{T},\\ \nabla\psi\cdot n(x)&=0\qquad\qquad\mbox{ for }(t,x)\in[0,T]\times\partial\Omega,\\ \psi(0,x)&=\psi_{0}(x)\qquad\mbox{ for }x\in\Omega.\end{aligned}\right.$
Then the following estimate holds
$\left\|\psi\right\|_{\mathrm{L}^{p}(\Omega_{T})}\leq\left(1+\Theta K_{\theta,\Theta,p^{\prime}}\right)T^{\frac{1}{p}}\left\|\psi_{0}\right\|_{\mathrm{L}^{p}(\Omega)},$
where the constant $K_{\theta,\Theta,p^{\prime}}$ is given by
$K_{\theta,\Theta,p^{\prime}}:=\frac{C^{PRC}_{\frac{\theta+\Theta}{2},p^{\prime}}\left(\frac{\Theta-\theta}{2}\right)}{1-C^{PRC}_{\frac{\theta+\Theta}{2},p^{\prime}}\left(\frac{\Theta-\theta}{2}\right)}\qquad\mbox{ provided we have }\qquad C^{PRC}_{\frac{\theta+\Theta}{2},p^{\prime}}\left(\frac{\Theta-\theta}{2}\right)<1.$
Here, the constant $C^{PRC}_{r,p^{\prime}}$ is the best constant in the following parabolic regularity estimate:
$\left\|\Delta\phi\right\|_{\mathrm{L}^{p^{\prime}}(\Omega_{T})}\leq C^{PRC}_{r,p^{\prime}}\left\|f\right\|_{\mathrm{L}^{p^{\prime}}(\Omega_{T})},$
where $\phi,f:[0,T]\times\Omega\to\mathbb{R}$ are any two functions such that $f\in\mathrm{L}^{p^{\prime}}(\Omega_{T})$ and they satisfy
$\partial_{t}\phi+r\Delta\phi&=f\qquad\mbox{ for }(t,x)\in\Omega_{T},\\ \nabla\phi\cdot n(x)&=0\qquad\mbox{ for }(t,x)\in[0,T]\times\partial\Omega,\\ \phi(T,x)&=0\qquad\mbox{ for }x\in\Omega.\end{aligned}\right.$
It has to be noted that $C^{PRC}_{r,p^{{}^{\prime}}}<\infty$ for $r>0$ and $C^{PRC}_{r,2}\leq\frac{1}{r}$ and depends only on $r,p^{{}^{\prime}}$, the domain and on the dimension, i.e it is independent of time. Moreover, as $C^{PRC}_{r,p^{{}^{\prime}}}<\infty$, if we take the difference between $\theta$ and $\Theta$ sufficiently small, then we have the required property that
$C^{PRC}_{\frac{\theta+\Theta}{2},p^{\prime}}\left(\frac{\Theta-\theta}{2}\right)<1.$
Proof of the above theorem can be found in [CDF14, Proposition 1.1].